### v1

# Can Nash inform capital requirements? Allocating systemic risk measures

Abstract
Systemic risk measures aggregate the risks from multiple financial institutions to find system-wide capital requirements. Though much attention has been given to assessing the level of systemic risk, less has been given to allocating that risk to the constituent institutions. Within this work, we propose a Nash allocation rule that is inspired by game theory. Intuitively, to construct these capital allocations, the banks compete in a game to reduce their own capital requirements while, simultaneously, maintaining system-level acceptability. We provide sufficient conditions for the existence and uniqueness of Nash allocation rules, and apply our results to the prominent structures used for systemic risk measures in the literature. We demonstrate the efficacy of Nash allocations with numerical case studies using the Eisenberg-Noe aggregation mechanism.
Key words: systemic risk, capital allocation, risk measure, game theory, Eisenberg-Noe model.
1 Introduction
The financial system is comprised of complex relationships and interconnections between different firms. Because of these connections, systemic risk naturally arises, i.e., when one firm or sector is shocked, negative externalities are felt by all other institutions. In its extreme, the distress or failure of one institution can lead to forms of financial contagion which amplify that initial shock throughout the whole of the financial system. While there are many models of systemic risk, e.g., for default contagion (Eisenberg and Noe, 2001; Rogers and Veraart, 2013) or price-mediated contagion (Cifuentes et al., 2005; Amini et al., 2016; Feinstein, 2017), within this work we take a model-agnostic view of systemic risk.
As systemic risk concerns the risk of the entire financial system, rather than ex ante focusing on any particular portfolio or firm, many metrics have been developed to assess such risks. Taking the idea from monetary risk measures (Artzner et al., 1999; Föllmer and Schied, 2002), systemic risk measures have been introduced that provide system-wide capital requirements to guarantee that the outcome to the entire system is acceptable. Within the literature, there are three broad notions of systemic risk measures:
  • Insensitive systemic risk measures: First introduced in Chen et al. (2013); Kromer et al. (2016), these systemic risk measures take the concept of aggregate-then-allocate. That is, the risks of the financial system are aggregated into a single random variable on which the system-wide capital requirements are determined.
  • Sensitive systemic risk measures: First introduced in Feinstein et al. (2017); Armenti et al. (2018), these systemic risk measures consider the converse of the insensitive systemic risk measures in that they allocate-then-aggregate. That is, each bank is assessed a capital charge which is included into the financial system before the aggregation to the system-wide risks is determined. Because there may be multiple Pareto-efficient allocations, these systemic risk measures are often presented as set-valued risk measures. Notably, and as highlighted in Feinstein et al. (2017), this approach includes seemingly disparate methodologies such as CoVaR (Adrian and Brunnermeier, 2016).
  • Scenario-dependent systemic risk measures: First introduced in Biagini et al. (2019), these systemic risk measures also consider the allocate-then-aggregate paradigm. In contrast to the sensitive systemic risk measures, the capital requirements are placed in a recovery fund to be distributed to any bank during a stress event so as to rescue the system. Because these recovery funds are distributed during the stress event, the distribution of capital is scenario-dependent.
In this way, systemic risk measures provide either the system-wide capital requirements (i.e., insensitive or scenario-dependent systemic risk measures) or sets of acceptable capital allocations (i.e., sensitive systemic risk measures). However, except in special cases highlighted in the below literature review (Section 1.1), no general methodology has been proposed to jointly select the capital requirements for all banks. We seek to develop a framework to construct capital requirements for all banks based on any systemic risk measure.
Within this work, we approach this problem with a game-theoretic construction for capital requirements. In utilizing game theory, we naturally capture the competition between banks trying to lower their own capital requirements while still maintaining systemic acceptability. Our goal is to systematically define these Nash allocation rules and determine sufficient conditions so as to guarantee their existence and uniqueness.
In systematizing the construction of systemic risk-aware capital allocations, we wish to highlight three primary contributions of this work. First, as far as the authors are aware, we are the first to propose a generic framework to allocate systemic risk capital requirements to the banks in the system. As we demonstrate within this work, the approach taken herein applies to all systemic risk measures that have been proposed in the prior literature. Notably, not only is this framework general, but due to its construction via a game, the proposed allocation mechanism is financially meaningful in all settings. Second, due to the game-theoretic construction, we prove general conditions for existence and uniqueness so that this construction can be guaranteed to function in practice. Third, we consider the Eisenberg and Noe (2001) aggregation function so as to demonstrate the proposed methodology when applied to financial networks. Within this structure, we provide a single convex optimization problem to compute a Nash allocation. This is in contrast to the traditional methodologies for systemic risk measures on financial networks that suffer from the curse of dimensionality (e.g., Feinstein et al. (2017)) or utilize approximation methods from, e.g., machine learning (Doldi et al., 2023; Gonon et al., 2024; Feinstein et al., 2023). We conclude with numerical case studies to demonstrate the efficacy of the Nash allocations when applied to financial networks.
The remainder of the main part of this work is organized as follows. The introduction is concluded with a literature review of other capital allocation rules in Section 1.1. Section 2 provides the general notation and definitions that are used throughout this work. The main results on the construction of a Nash allocation rule, including sufficient conditions for the existence and uniqueness of that rule, are provided in Section 3. Nash allocation rules are further studied in detail and demonstrated in numerical case studies in Section 4 on systemic risk measures generated by the Eisenberg and Noe (2001) clearing system. Section 5 concludes the work.
1.1 Literature Review and Systemic Capital Allocations
Though the aforementioned systemic risk measures are based on the idea of capital requirements, there has been only limited consideration for how to determine the actual capital requirement to be charged to each institution. The first attempt at determining the capital allocations for sensitive systemic risk measures was proposed in Feinstein et al. (2017), and independently developed in Armenti et al. (2018), via the efficient cash-invariant allocation rules. These allocations are determined by minimizing the total system-wide capital required and they are studied further in, e.g., Fissler et al. (2021). In contrast, Brunnermeier and Cheridito (2019) presents economic axioms for a capital allocation; they present a construction for such a capital allocation in the setting of a specific sensitive systemic risk measure.
There have been different approaches taken for determining the allocation of the individual capital requirements for the scenario-dependent systemic risk measures. Within Biagini et al. (2020), these allocations are determined via the dual representation of the systemic risk measure where the system health is determined by the sum of bank utilities. A separate allocation approach is presented in Wang et al. (2023) in which multiple optimization problems need to be solved to determine the allocations.
A different approach, in which banks are allowed to directly exchange their risks, is taken in Biagini et al. (2021); Doldi and Frittelli (2024). However, in doing so, the authors did not directly consider the question of capital requirement but rather the optimal sharing of an exogenously provided rescue fund. Though we will directly consider the question of capital requirements, we note that the approach taken herein is most similar to this risk-transfer approach in that we will consider how banks may interact with each other to reduce their capital requirements.
Finally, due to the use of risk measures to compute systemic capital requirements, it may be tempting to apply a vector-valued risk measure to provide individualized capital requirements while maintaining the interconnections between firms. However, as proven in Ararat and Feinstein (2024), no lower semicontinuous, convex vector-valued risk measure can admit capital requirements that include systemic interactions. That is, following a vector-valued risk measure, the capital requirements for every bank can only depend on its own financial situation. In this way, such a methodology is incapable of capturing the interconnections that are inherent in financial systems.
2 Background and Notation
Before discussing the relevant background on risk measures, we provide some overarching notation that is used throughout this work. Let us consider the Euclidean space $\mathbb{R}^{N}$, where $N\in\mathbb{N}:=\{1,2,\ldots\}$. For each $i\in[N]\coloneqq\{1,\ldots,N\}$, we denote by $e_{i}$ the $i^{\text{th}}$ standard unit vector in $\mathbb{R}^{N}$. Given $x=(x_{1},\ldots,x_{N})^{\top}\in\mathbb{R}^{N}$, $m\in\mathbb{R}$, $i\in[N]$, we define $(m,x_{-i})\in\mathbb{R}^{N}$ to be the vector whose $i^{\text{th}}$ component is $m$ and who is equal to $x$ in every other component. To compare two vectors $x,y\in\mathbb{R}^{N}$ componentwise, we write $x\leq y$ if $x_{i}\leq y_{i}$ for each $i\in[N]$. In this case, we also define the hyperrectangle $[x,y]:=\{z\in\mathbb{R}^{N}\;|\;x\leq z\leq y\}$. We say that a function $f\colon\mathbb{R}^{N}\to\bar{\mathbb{R}}\coloneqq[-\infty,+\infty]$ is increasing if $x\leq y$ implies $f(x)\leq f(y)$ for every $x,y\in\mathbb{R}^{N}$; we say that $f$ is decreasing if $-f$ is increasing. Since we will mostly deal with concave aggregation functions in the sequel, we define the effective domain of a function $f\colon\mathbb{R}^{N}\to\bar{\mathbb{R}}$ as
$\operatorname{dom}f\coloneqq\{x\in\mathbb{R}^{N}\;|\;f(x)>-\infty\}.$
For a sequence $(x^{n})_{n\in\mathbb{N}}$ and a number $x$ in $\bar{\mathbb{R}}$, we write $x^{n}\uparrow x$ ($x^{n}\downarrow x$) if $x^{1}\leq x^{2}\leq\ldots$ ($x^{1}\geq x^{2}\geq\ldots$) and $\lim_{n\rightarrow\infty}x^{n}=x$. We denote by $I_{N}$ the $N\times N$ identity matrix. For a matrix $a\in\mathbb{R}^{N\times M}$, where $M\in\mathbb{N}$, we denote by $a_{i\cdot}$ its $i^{\text{th}}$ row (treated as a column vector) and by $a_{\cdot j}$ its $j^{\text{th}}$ column for each $i\in[N]$ and $j\in[M]$. For sets $A,B\subseteq\mathbb{R}^{N}$, we define their Minkowski sum as $A+B:=\{x+y\;|\;x\in A,y\in B\}$ with the convention $A+\emptyset=\emptyset+B=\emptyset$. Then, the collection of all upper sets in $\mathbb{R}^{N}$ is given by
$\mathcal{P}(\mathbb{R}^{N};\mathbb{R}^{N}_{+})\coloneqq\{A\subseteq\mathbb{R}^{N}\;|\;A+\mathbb{R}^{N}_{+}=A\},$ (2.1)
[…]
${}.\mathcal{R}(\mathbb{R}^{N};\mathbb{R}^{N}_{+})\coloneqq\left\{\bigtimes_{i=1}^{N}A_{i}\mid\forall i\in[N]\colon A_{i}=\mathbb{R}\text{ or }A_{i}=[x_{i},+\infty)\text{ for some }x_{i}\in\mathbb{R}\right\}\cup\{\emptyset\}.$ (2.2)
Clearly, $\mathcal{R}(\mathbb{R}^{N};\mathbb{R}^{N}_{+})\subseteq\mathcal{P}(\mathbb{R}^{N};\mathbb{R}^{N}_{+})$.
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space. We denote the space of all (equivalence classes of) $\mathcal{F}$-measurable $N$-dimensional random vectors $X=(X_{1},\ldots,X_{N})^{\top}$ by $L^{0}(\mathbb{R}^{N})$. In particular, (in)equalities and limits about random variables are understood in the $\mathbb{P}$-almost sure (a.s.) sense unless stated otherwise. For a random variable $Y\in L^{0}(\mathbb{R})$, we define its $\mathbb{P}$-essential supremum as $\operatorname{ess\,sup}Y:=\inf\{c>0\;|\;\mathbb{P}\{Y\leq c\}=1\}$ and $\mathbb{P}$-essential infimum as $\operatorname{ess\,inf}Y:=-\operatorname{ess\,sup}(-Y)$. For mathematical convenience, we often restrict ourselves to the space $L^{\infty}(\mathbb{R}^{N})\subseteq L^{0}(\mathbb{R}^{N})$ of essentially bounded random vectors, i.e., the set of all $X\in L^{0}(\mathbb{R}^{N})$ such that $\|X_{i}\|_{\infty}:=\max\{\operatorname{ess\,sup}X_{i},-\operatorname{ess\,inf}X_{i}\}<+\infty$ for every $i\in[N]$. With abuse of notation, we often write $\operatorname{ess\,sup}X:=(\operatorname{ess\,sup}X_{1},\ldots,\operatorname{ess\,sup}X_{N})^{\top}\in\mathbb{R}^{N}$ and $\operatorname{ess\,inf}X:=(\operatorname{ess\,inf}X_{1},\ldots,\operatorname{ess\,inf}X_{N})^{\top}\in\mathbb{R}^{N}$ for $X\in L^{\infty}(\mathbb{R}^{N})$.
2.1 Monetary Risk Measures
Let us recall the definition of a risk measure in the univariate setting.
Definition 2.1.
A functional $\rho\colon L^{\infty}(\mathbb{R})\to\mathbb{R}$ is called a risk measure if it satisfies the following properties:
  • Normalized: $\rho(0)=0$.
  • Monotone: $Y^{1}\geq Y^{2}$ implies $\rho(Y^{1})\leq\rho(Y^{2})$ for every $Y^{1},Y^{2}\in L^{\infty}(\mathbb{R})$.
  • Translative: $\rho(Y+m)=\rho(Y)-m$ for every $Y\in L^{\infty}(\mathbb{R})$ and $m\in\mathbb{R}$.
In addition, we usually consider the following properties for $\rho$:
  • Convex: $\rho(\lambda Y^{1}+(1-\lambda)Y^{2})\leq\lambda\rho(Y^{1})+(1-\lambda)\rho(Y^{2})$ for every $Y^{1},Y^{2}\in L^{\infty}(\mathbb{R})$ and $\lambda\in[0,1]$.
  • Positively homogeneous: $\rho(\lambda Y)=\lambda\rho(Y)$ for every $Y\in L^{\infty}(\mathbb{R})$ and $\lambda\geq 0$.
  • Continuous from above: $Y^{n}\downarrow Y$ implies $\rho(Y^{n})\uparrow\rho(Y)$ for every $Y,Y^{1},Y^{2},\ldots\in L^{\infty}(\mathbb{R})$.
A risk measure $\rho$ is called coherent if it is convex and positively homogeneous.
Given a risk measure $\rho$, the set
$\mathcal{A}:=\{Y\in L^{\infty}(\mathbb{R})\;|\;\rho(Y)\leq 0\}$
is called its acceptance set. As recalled in the next proposition, this set characterizes $\rho$.
Proposition 2.2.
(Föllmer and Schied, 2016, Proposition 4.6) Let $\rho\colon L^{\infty}(\mathbb{R})\to\mathbb{R}$ be a risk measure with acceptance set $\mathcal{A}$. Then, for every $Y\in L^{\infty}(\mathbb{R})$, it holds
$\rho(Y)=\inf\{m\in\mathbb{R}\;|\;Y+m\in\mathcal{A}\}.$
Moreover, $\rho$ is convex (positively homogeneous, coherent) if and only if $\mathcal{A}$ is a convex set (cone, convex cone); whenever $\rho$ is convex, $\rho$ is continuous from above if and only if $\mathcal{A}$ is closed in the weak* topology on $L^{\infty}(\mathbb{R})$.
2.2 Systemic Risk Measures
Consider a financial system with $N\in\mathbb{N}$ institutions. In this multivariate setting, we will consider systemic risk measures that are defined as set-valued functionals on the space $L^{\infty}(\mathbb{R}^{N})$. Here, each $X\in L^{\infty}(\mathbb{R}^{N})$ is called a random shock or a stress scenario, e.g., for each $i\in[N]$, the random variable $X_{i}$ represents the future worth of the assets of institution $i$, which affects its ability to meet its contractual obligations.
We start by introducing the notion of aggregation function in our setting. For a function $\Lambda\colon\mathbb{R}^{N}\times\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ and $\bar{x},\bar{m}\in\mathbb{R}^{N}$, we denote by $\Lambda|_{m=\bar{m}}$ the section $x\mapsto\Lambda(x,\bar{m})$ and by $\Lambda|_{x=\bar{x}}$ the section $m\mapsto\Lambda(\bar{x},m)$. Recall that we consider the effective domain $\operatorname{dom}\Lambda=\{(x,m)\in\mathbb{R}^{N}\times\mathbb{R}^{N}\;|\;\Lambda(x,m)\in\mathbb{R}\}$; in particular, $\Lambda(x,m)=-\infty$ for $(x,m)\not\in\operatorname{dom}\Lambda$.
Definition 2.3.
A function $\Lambda\colon\mathbb{R}^{N}\times\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ is called an aggregation function if it satisfies the following properties:
  1. 1.
    Monotone: $\Lambda$ is increasing with $\operatorname{dom}\Lambda\neq\emptyset$.
  2. 2.
    Upper semicontinuous: $\Lambda$ is upper semicontinuous.
  3. 3.
    Rectangular domain: (i) For each $\bar{m}\in\mathbb{R}^{N}$, it holds $\operatorname{dom}\Lambda|_{m=\bar{m}}\in\mathcal{R}(\mathbb{R}^{N};\mathbb{R}^{N}_{+})$.
    (ii) For each $\bar{x}\in\mathbb{R}^{N}$, it holds $\operatorname{dom}\Lambda|_{x=\bar{x}}\in\mathcal{R}(\mathbb{R}^{N};\mathbb{R}^{N}_{+})$.
A function $\bar{\Lambda}\colon\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ is called a single-element aggregation function if $(x,m)\mapsto\bar{\Lambda}(x)$ is an aggregation function.
Remark 2.4.
In view of Definition 2.3, a function $\bar{\Lambda}\colon\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ is a single-element aggregation function if and only if it is increasing, upper semicontinuous, and $\operatorname{dom}\bar{\Lambda}$ is a nonempty rectangular upper set. In particular, part (ii) of the rectangular domain property becomes redundant in this case.
Remark 2.5.
For an aggregation function $\Lambda$, the quantity $\Lambda(x,m)$ can be seen as the impact of a shock of magnitude $x\in\mathbb{R}^{N}$ to society under capital level $m\in\mathbb{R}^{N}$. More generally, we can easily extend Definition 2.3 to functions $\Lambda\colon\mathbb{R}^{N}\times\mathbb{R}^{M}\to\mathbb{R}\cup\{-\infty\}$, where $M\in\mathbb{N}$. In this way, the shocks need not be of the same dimension as the capital requirements. This applies if, e.g., we consider factor models to generate the shocks (see Banerjee and Feinstein (2022)) or groupings of the institutions for capital requirements (see Feinstein et al. (2017); Ararat and Meimanjan (2023)).
We will discuss single-element aggregation functions later starting at Example 2.9. For our general framework, let $\mathcal{A}\subseteq L^{\infty}(\mathbb{R})$ be the acceptance set of a risk measure $\rho$ and $\Lambda\colon\mathbb{R}^{N}\times\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ be an aggregation function. For each $X\in L^{\infty}(\mathbb{R}^{N})$, we define the set
$D(X)\coloneqq\{m\in\mathbb{R}^{N}\;|\;\mathbb{P}\{(X,m)\in\operatorname{dom}\Lambda\}=1\},$
i.e., $\Lambda(X,m)$ is a real-valued random variable for each $m\in D(X)$. In the sequel, it will be convenient to introduce the set
$L^{\infty}_{\Lambda}(\mathbb{R}^{N})\coloneqq\left\{X\in L^{\infty}(\mathbb{R}^{N})\;|\;D(X)\neq\emptyset\right\}.$
Since $\operatorname{dom}\Lambda\neq\emptyset$, it is easy to verify that $L^{\infty}_{\Lambda}(\mathbb{R}^{N})\neq\emptyset$ as well.
Lemma 2.6.
(i) For each $X\in L^{\infty}(\mathbb{R}^{N})$, we have $D(X)=D(\operatorname{ess\,inf}X)=\operatorname{dom}\Lambda|_{x=\operatorname{ess\,inf}X}$.
(ii) For each $X\in L_{\Lambda}^{\infty}(\mathbb{R}^{N})$ and $m\in D(X)$, it holds $\Lambda(X,m)\in L^{\infty}(\mathbb{R})$.
Proof.
(i) Let $X\in L^{\infty}(\mathbb{R}^{N})$. Since $\operatorname{ess\,inf}X\leq X$ a.s., we have $D(\operatorname{ess\,inf}X)\subseteq D(X)$. The converse is trivial if $D(X)=\emptyset$. Otherwise, let $\bar{m}\in D(X)$. Then, $\mathbb{P}\{X\in\operatorname{dom}\Lambda|_{m=\bar{m}}\}=1$. In particular, $\operatorname{dom}\Lambda|_{m=\bar{m}}$ is a nonempty set. Since $\operatorname{dom}\Lambda|_{m=\bar{m}}\in\mathcal{R}(\mathbb{R}^{N}_{+};\mathbb{R}^{N}_{+})$, there exists $\underline{x}\in\operatorname{dom}\Lambda|_{m=\bar{m}}$ such that $\underline{x}\leq X$ a.s. Then, $\underline{x}\leq\operatorname{ess\,inf}X$ so that $-\infty<\Lambda(\underline{x},\bar{m})\leq\Lambda(\operatorname{ess\,inf}X,\bar{m})$ by the monotonicity of $\Lambda$. Hence, $\bar{m}\in D(\operatorname{ess\,inf}X)$ and we get $D(X)\subseteq D(\operatorname{ess\,inf}X)$.
(ii) Let $X\in L_{\Lambda}^{\infty}(\mathbb{R}^{N})$ and $m\in D(X)$. By (i), we have $m\in D(\operatorname{ess\,inf}X)$ so that $-\infty<\Lambda(\operatorname{ess\,inf}X,m)\leq\Lambda(X,m)\leq\Lambda(\operatorname{ess\,sup}X,m)$ a.s. by monotonicity. Hence, we obtain $\Lambda(X,m)\in L^{\infty}(\mathbb{R})$. ∎
Definition 2.7.
The set-valued functional $R\colon L^{\infty}_{\Lambda}(\mathbb{R}^{N})\rightrightarrows\mathbb{R}^{N}$ defined by
$R(X)\coloneqq\{m\in D(X)\;|\;\Lambda(X,m)\in\mathcal{A}\},$
for each $X\in L_{\Lambda}^{\infty}(\mathbb{R}^{N})$, is called the $(\mathcal{A},\Lambda)$-systemic risk measure.
In the above definition, if a stress scenario $X\in L_{\Lambda}^{\infty}(\mathbb{R}^{N})$ is given, then $R(X)$ provides the set of all acceptable systemic capital allocations, i.e., accounting for interactions between institutions as modeled by the aggregation function $\Lambda$.
Remark 2.8.
Since $m\mapsto\Lambda(x,m)$ is increasing with a rectangular domain for each $x\in\mathbb{R}^{N}$, the set-valued function $D$ and the $(\mathcal{A},\Lambda)$-systemic risk measure $R$ map into the collection of rectangular upper and upper sets, respectively, i.e., $D(X)\in\mathcal{R}(\mathbb{R}^{N};\mathbb{R}^{N}_{+})$ and $R(X)\in\mathcal{P}(\mathbb{R}^{N};\mathbb{R}^{N}_{+})$ for each $X\in L_{\Lambda}^{\infty}(\mathbb{R}^{N})$; see (2.1) and (2.2). Moreover, since $x\mapsto\Lambda(x,m)$ is increasing for each $m\in\mathbb{R}^{N}$ as well, $D$ and $R$ have the following monotonicity property: $D(X^{1})\subseteq D(X^{2})$ and $R(X^{1})\subseteq R(X^{2})$ for every $X^{1},X^{2}\in L^{\infty}_{\Lambda}(\mathbb{R}^{N})$ with $X^{1}\leq X^{2}$.
We proceed with examples of systemic risk measures that have been studied in the recent literature.
Example 2.9.
Within the literature, systemic risk measures tend to follow one of three forms discussed below. All of these constructions are derived from a single-element aggregation function $\bar{\Lambda}\colon\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$.
  • Insensitive systemic risk measures: Let us take $\Lambda=\Lambda^{I}$, where
    $\Lambda^{I}(x,m):=\bar{\Lambda}(x)+\sum_{i\in[N]}m_{i}$ (2.3)
    for each $x,m\in\mathbb{R}^{N}$. Clearly, $\Lambda^{I}$ is an aggregation function and $\operatorname{dom}\Lambda^{I}=\operatorname{dom}\bar{\Lambda}\times\mathbb{R}^{N}$. Let
    $L^{\infty}(\operatorname{dom}\bar{\Lambda})\coloneqq\{X\in L^{\infty}(\mathbb{R}^{N})\;|\;\mathbb{P}\{X\in\operatorname{dom}\bar{\Lambda}\}=1\}.$
    Then, we have $D^{I}(X):=D(X)=\emptyset$ if $X\notin L^{\infty}(\operatorname{dom}\bar{\Lambda})$ and $D^{I}(X):=D(X)=\mathbb{R}^{N}$ if $X\in L^{\infty}(\operatorname{dom}\bar{\Lambda})$. In particular, $L^{\infty}_{\Lambda^{I}}(\mathbb{R}^{N})=L^{\infty}(\operatorname{dom}\bar{\Lambda})$. For every $X\in L^{\infty}(\operatorname{dom}\bar{\Lambda})$, by Lemma 2.6, we have $\bar{\Lambda}(X)\in L^{\infty}(\mathbb{R})$ and
    $R^{I}(X)\coloneqq R(X)\negthinspace=\negthinspace\left\{m\in\mathbb{R}^{N}\;|\;\bar{\Lambda}(X)+\sum_{i\in[N]}m_{i}\in\mathcal{A}\right\}\negthinspace=\negthinspace\left\{m\in\mathbb{R}^{N}\;|\;\rho(\bar{\Lambda}(X))\leq\sum_{i\in[N]}m_{i}\right\};$
    in particular, $R^{I}(X)$ is a halfspace. Hence, the set-valued functional $R^{I}$ essentially reduces to the real-valued functional $X\mapsto\rho(\bar{\Lambda}(X))$ on $L^{\infty}(\operatorname{dom}\bar{\Lambda})$ as studied in Chen et al. (2013); Kromer et al. (2016).
  • Sensitive systemic risk measures: Let us take $\Lambda=\Lambda^{S}$, where
    $\Lambda^{S}(x,m)\coloneqq\bar{\Lambda}(x+m)$ (2.4)
    for each $x,m\in\mathbb{R}^{N}$. Clearly, $\Lambda^{S}$ is an aggregation function with
    $\operatorname{dom}\Lambda^{S}=\{(x,m)\in\mathbb{R}^{N}\times\mathbb{R}^{N}\;|\;x+m\in\operatorname{dom}\bar{\Lambda}\}.$ (2.5)
    Then, for each $X\in L^{\infty}(\mathbb{R}^{N})$, we have
    $D^{S}(X):=D(X)=\{m\in\mathbb{R}^{N}\;|\;\mathbb{P}\{X+m\in\operatorname{dom}\bar{\Lambda}\}=1\}.$ (2.6)
    Since $\operatorname{dom}\bar{\Lambda}\neq\emptyset$ and $\bar{\Lambda}$ is increasing, it follows that $D^{S}(X)\neq\emptyset$ for every $X\in L^{\infty}(\mathbb{R}^{N})$. In particular, $L^{\infty}_{\Lambda^{S}}(\mathbb{R}^{N})=L^{\infty}(\mathbb{R}^{N})$. By Lemma 2.6, we have $\bar{\Lambda}(X+m)\in L^{\infty}(\mathbb{R})$ for every $X\in L^{\infty}(\mathbb{R}^{N})$ and $m\in D^{S}(X)$. Then,
    $R^{S}(X)\coloneqq R(X)$ =$=\{m\in D^{S}(X)\;|\;\bar{\Lambda}(X+m)\in\mathcal{A}\}$
    =$=\{m\in\mathbb{R}^{N}\;|\;\rho(\bar{\Lambda}(X+m))\leq 0,\ \operatorname{ess\,inf}X+m\in\operatorname{dom}\bar{\Lambda}\}$
    for every $X\in L^{\infty}(\mathbb{R}^{N})$, which yields the systemic risk measures studied in Feinstein et al. (2017); Ararat and Rudloff (2020). Compared to the previous case, these systemic risk measures consider the direct impact of capital allocations on the stress scenario before aggregation, hence they are more sensitive to capital levels.
  • Scenario-dependent systemic risk measures: Suppose that $\operatorname{dom}\bar{\Lambda}=\underline{x}+\mathbb{R}^{N}_{+}$ for some $\underline{x}\in\mathbb{R}^{N}$. Let us take $\Lambda=\Lambda^{SD}$, where
    $\Lambda^{SD}(x,m)$$\coloneqq\begin{cases}\sup_{y\in\mathbb{R}^{N}}\{\bar{\Lambda}(x+y)\;|\;\sum_{i\in[N]}y_{i}\leq\sum_{i\in[N]}m_{i}\}&\text{if }x+m\in\operatorname{dom}\bar{\Lambda}\\ -\infty&\text{else}\end{cases}$ (2.7)
    for each $x,m\in\mathbb{R}^{N}$. Note that, when $x+m\in\operatorname{dom}\bar{\Lambda}$, the feasible region of the maximization problem in (2.7) is $\{y\geq\underline{x}-x\;|\;\sum_{i\in[N]}y_{i}\leq\sum_{i\in[N]}m_{i}\}$, which is a nonempty, compact set, and we have $\Lambda^{SD}(x,m)\in\mathbb{R}$. In particular,
    $\operatorname{dom}\Lambda^{SD}=\operatorname{dom}\Lambda^{S},\quad D^{SD}(X):=D(X)=D^{S}(X),\quad L^{\infty}_{\Lambda^{SD}}(\mathbb{R}^{N})=L^{\infty}(\mathbb{R}^{N});$
    see (2.5) and (2.6). Since $\bar{\Lambda}$ is upper semicontinuous, it follows from (Rockafellar and Wets, 2009, Theorem 1.17) that $\Lambda^{SD}$ is upper semicontinuous as well. Hence, $\Lambda^{SD}$ is an aggregation function. Then, for every $X\in L^{\infty}(\mathbb{R}^{N})$, it is easy to verify that
    $R^{SD}(X)$ $\coloneqq R(X)=\{m\in D^{SD}(X)\;|\;\Lambda^{SD}(X,m)\in\mathcal{A}\}$
[…]
    for every $X\in L^{\infty}(\mathbb{R}^{N})$. In particular, $R^{SD}(X)$ is the intersection of a halfspace with $D^{SD}(X)$ for every $X\in L^{\infty}(\mathbb{R}^{N})$. Hence, the set-valued functional $R^{SD}$ essentially reduces to the minimum sum scalarization $X\mapsto\inf\{\sum_{i\in[N]}m_{i}\;|\;m\in R^{SD}(X)\}$ as studied in Biagini et al. (2019, 2020, 2021).
We conclude our review of systemic risk measures by providing examples of single-element aggregation functions that we will revisit in Section 3 below. We note that we also consider a different aggregation function based on financial clearing conditions in Section 4.
Example 2.10.
Here are a few examples of single-element aggregation functions:
  1. 1.
    Sum of univariate utilities: For each $i\in[N]$, let $u_{i}\colon\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$ be a univariate utility function, i.e., an increasing upper semicontinuous function such that $\operatorname{dom}u_{i}\neq\emptyset$. Let $\bar{\Lambda}^{U}(x)\coloneqq\sum_{i\in[N]}u_{i}(x_{i})$ for each $x\in\mathbb{R}^{N}$. As presented in, e.g., Biagini et al. (2019, 2020), the function $\bar{\Lambda}^{U}$ is a single-element aggregation function with $\operatorname{dom}\bar{\Lambda}^{U}=\bigtimes_{i\in[N]}\operatorname{dom}u_{i}\neq\emptyset$.
  2. 2.
    Fund of funds: As detailed in Chen et al. (2013, Example 3), when an individual invests with a hedge fund, they are charged a fee on their gains. That is, a gain of $x_{i}\in\mathbb{R}$ in firm $i\in[N]$ has a payout of $x_{i}-\gamma_{i}x_{i}^{+}$, where $\gamma_{i}\in[0,1)$ is the fee rate for firm $i$. However, it is possible for the investor to not just invest in individual firms, but in a fund of funds, where an additional $\gamma\in[0,1)$ fraction of the profits of that aggregate fund is taken by the fund administrator. From this construction, we define
    $\bar{\Lambda}^{FoF}(x):=\sum_{i\in[N]}(x_{i}-\gamma_{i}x_{i}^{+})-\gamma\left(\sum_{i\in[N]}(x_{i}-\gamma_{i}x_{i}^{+})\right)^{+}$
    for each $x\in\mathbb{R}^{N}$. Clearly, $\bar{\Lambda}^{FoF}$ is a continuous single-element aggregation function with $\operatorname{dom}\bar{\Lambda}^{FoF}=\mathbb{R}^{N}$.
  3. 3.
    Mean-field utility: Inspired by mean-field games (see, e.g., Carmona et al. (2013, Section 3.1) and Bensoussan et al. (2016, Section 2)), we introduce the idea that each bank’s utility depends on itself and the average health of the financial system. Mathematically, each bank $i\in[N]$ has utility $u_{i}(x_{i},\frac{1}{N}\sum_{j\in[N]}x_{j})$ so that the utility of bank $i$ depends on itself ($x_{i}$) and an aggregate of the financial system ($\frac{1}{N}\sum_{j\in[N]}x_{j}$), where $u_{i}\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$ is an aggregation function with $\operatorname{dom}u_{i}=A_{i}\times\bar{A}$ for some $A_{i}\in\mathcal{R}(\mathbb{R};\mathbb{R}_{+})\setminus\{\emptyset\}$ (depending on $i$) and $\bar{A}\in\mathcal{R}(\mathbb{R};\mathbb{R}_{+})\setminus\{\emptyset\}$ (independent of $i$). We also assume that ×$\bigtimes_{i\in[N]}A_{i}\subseteq\{x\in\mathbb{R}^{N}\;|\;\frac{1}{N}\sum_{j\in[N]}x_{j}\in\bar{A}\}$. As in the case of sum of univariate utilities, we define
    $\bar{\Lambda}^{MF}(x):=\sum_{i\in[N]}u_{i}\left(x_{i},\frac{1}{N}\sum_{j\in[N]}x_{j}\right)$
    for each $x\in\mathbb{R}^{N}$, which is a single-element aggregation function with $\operatorname{dom}\bar{\Lambda}^{MF}=\bigtimes_{i\in[N]}A_{i}$.
3 Nash Allocation Rules
In this section, we wish to construct a financially meaningful capital allocation rule that accounts for the contribution that each firm makes towards the risk of the full system. We are specifically motivated by the aggregation functions that are sums of univariate utilities (see Example 2.10(1)), i.e., $\bar{\Lambda}(x)=\sum_{i\in[N]}u_{i}(x_{i})$, $x\in\mathbb{R}^{N}$, for univariate utility functions $u_{1},\ldots,u_{N}\colon\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$ so that each bank’s contribution to the system’s utility is clear. We aim to generalize this concept for broader classes of aggregation functions. Notably, in contrast to the set-valued systemic risk measures introduced in Section 2.2, which provide the collection of all acceptable capital requirements, we focus on finding specific, i.e., vector-valued, acceptable capital allocations instead.
Throughout this section, let $\mathcal{A}$ be the acceptance set of a risk measure $\rho$ and $\Lambda\colon\mathbb{R}^{N}\times\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ be an aggregation function. We consider the corresponding $(\mathcal{A},\Lambda)$-systemic risk measure $R$ given by Definition 2.7. We will work under the following assumption:
Assumption 3.1.
The risk measure $\rho$ is a coherent risk measure that is continuous from above; equivalently, the acceptance set $\mathcal{A}$ is a convex cone that is closed with respect to the weak* topology on $L^{\infty}(\mathbb{R})$.
As a generalization of aggregation functions that are in the form of a sum of univariate utility functions, we introduce the notion of decomposition of an aggregation function next.
Definition 3.2.
Let $\Lambda_{1},\ldots,\Lambda_{N}\colon\mathbb{R}^{N}\times\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ be aggregation functions. The collection $(\Lambda_{i})_{i\in[N]}$ is called a decomposition of $\Lambda$ if the following properties hold:
  1. 1.
    Consistent domain: For each $i\in[N]$, it holds $\operatorname{dom}\Lambda_{i}=\operatorname{dom}\Lambda$.
  2. 2.
    Dominated by $\Lambda$: It holds $\sum_{i\in[N]}\Lambda_{i}(x,m)\leq\Lambda(x,m)$ for every $x,m\in\mathbb{R}^{N}$.
  3. 3.
    Self-feasible: For every $\bar{x}\in\mathbb{R}^{N}$ with $D(\bar{x})=\operatorname{dom}\Lambda|_{x=\bar{x}}\neq\emptyset$, there exists $\bar{r}\in D(\bar{x})$ such that $\Lambda_{i}(\bar{x},(\bar{r}_{i},m_{-i}))\geq 0$ for every $m\in D(\bar{x})$ and $i\in[N]$.
  4. 4.
    Self-infeasible: For every $\underline{x}\in\mathbb{R}^{N}$ with $D(\underline{x})=\operatorname{dom}\Lambda|_{x=\underline{x}}\neq\emptyset$, there exists $\underline{r}\in D(\underline{x})$ such that $-\infty\leq\Lambda_{i}(\underline{x},(\underline{r}_{i}-\varepsilon,m_{-i}))<0$ for every $m\in D(\underline{x})$, $i\in[N]$, and $\varepsilon>0$.
The decomposition $(\Lambda_{i})_{i\in[N]}$ is called self-preferential if there exists some constant $L\in[0,1)$ such that, for every $i,j\in[N]$ with $i\neq j$, $(x,m)\in\operatorname{dom}\Lambda$, and $\delta\geq 0$, it holds
$\Lambda_{i}(x,m+L\delta e_{i})\geq\Lambda_{i}(x,m+\delta e_{j}).$
Remark 3.3.
The self-preferential property of a decomposition means that increasing the capital to the bank associated with the aggregation function is more efficient (by a fixed factor) than adding it to another bank. In particular, from the perspective of a Nash game, no bank has an incentive to bail out any other bank; instead, banks will retain any capital they are allocated. This idea will be formally introduced in Definition 3.5 below.
Next, we illustrate Definition 3.2 through the motivating example of sum of univariate utilities aggregation. Decompositions for single-element aggregation functions will be defined and studied systematically in Section 3.3.
Example 3.4.
Recall the single-element aggregation function $\bar{\Lambda}^{U}$ in Example 2.10(1) that is in the form of sum of univariate utility functions $u_{1},\ldots,u_{N}$. For this example, consider the corresponding sensitive aggregation function $\Lambda^{S}$ defined by
$\Lambda^{S}(x,m)\coloneqq\bar{\Lambda}^{U}(x+m)=\sum_{i\in[N]}u_{i}(x_{i}+m_{i})$
for each $x,m\in\mathbb{R}^{N}$. For each $i\in[N]$, let us define
$\Lambda^{S}_{i}(x,m)\coloneqq u_{i}(x_{i}+m_{i})$
if $x,m\in\mathbb{R}^{N}$ satisfy $x+m\in\operatorname{dom}\bar{\Lambda}^{U}$ and $\Lambda^{S}_{i}(x,m)\coloneqq-\infty$ otherwise. Clearly, $\Lambda^{S}_{1},\ldots,\Lambda^{S}_{N}$ are aggregation functions, $(\Lambda^{S}_{i})_{i\in[N]}$ satisfies the consistent domain property, and $\sum_{i\in[N]}\Lambda^{S}_{i}=\Lambda^{S}$. Let us further assume that $\{x_{i}\in\operatorname{dom}u_{i}\;|\;u_{i}(x_{i})\geq 0\}\not\in\{\emptyset,\mathbb{R}\}$ for each $i\in[N]$, which guarantees self-feasibility and self-infeasibility for $(\Lambda^{S}_{i})_{i\in[N]}$. Hence, $(\Lambda^{S}_{i})_{i\in[N]}$ is a decomposition of $\Lambda^{S}$. Moreover, this decomposition is self-preferential with $L=0$ due to the independence of the utility functions.
Using a decomposition of $\Lambda$, we are ready to define a Nash allocation rule, the central object of the paper.
Definition 3.5.
Let $(\Lambda_{i})_{i\in[N]}$ be a decomposition of $\Lambda$. A vector-valued functional $r\colon L_{\Lambda}^{\infty}(\mathbb{R}^{N})\to\mathbb{R}^{N}$ is called a $(\mathcal{A},(\Lambda_{i})_{i\in[N]})$-Nash allocation rule if, for every $X\in L_{\Lambda}^{\infty}(\mathbb{R}^{N})$, it holds
$r_{i}(X)=\inf\{m_{i}\in\mathbb{R}\;|\;\Lambda_{i}(X,(m_{i},r_{-i}(X)))\in\mathcal{A}\}$
for each $i\in[N]$. [^1] In the infimum, we implicitly consider $m_{i}\in\mathbb{R}$ such that $(m_{i},r_{-i}(X))\in D(X)$, which ensures that $\Lambda_{i}(X,(m_{i},r_{-i}(X)))\in L^{\infty}(\mathbb{R})$ by Lemma 2.6. In this case, $r$ is called acceptable if $r(X)\in R(X)$ for every $X\in L_{\Lambda}^{\infty}(\mathbb{R}^{N})$.
In Appendix A, we propose a generalization of a Nash allocation rule that permits banks to use their capital to directly bail out or rescue other institutions. However, for all examples considered herein, all Nash allocations with rescue funds satisfy the above definition (see Corollary A.2); hence, we focus on this simpler setting in this paper. We leave a more formal study of this generalization for future work.
Within this section, we first consider general conditions for the acceptability of a Nash allocation rule (Section 3.1). With conditions so that acceptability is guaranteed, we provide a general theorem for the existence and uniqueness of Nash allocation rules (Section 3.2). With these results for general systemic risk measures, we provide further insights for those systemic risk measures generated from a single-element aggregation function, i.e., insensitive (Section 3.3.1), sensitive (Section 3.3.2), and scenario-dependent (Section 3.3.3) systemic risk measures.
3.1 Acceptable Allocation Rules
Before focusing on Nash allocation rules specifically, we want to consider the acceptability of a given capital allocation vector. Specifically, when we decompose an aggregation function into attributed pieces for each institution, it is not guaranteed that the acceptability of all banks will lead to systemic acceptability; we refer the interested reader to Example 3.7 for a simple example without systemic acceptability. Proposition 3.6 provides a simple sufficient condition for acceptability. Recall that Assumption 3.1 is imposed herein and for the remainder of this work.
Proposition 3.6.
Let $(\Lambda_{i})_{i\in[N]}$ be a decomposition of $\Lambda$. Consider a random shock $X\in L_{\Lambda}^{\infty}(\mathbb{R}^{N})$ and a capital vector $m\in D(X)$. If $\Lambda_{i}(X,m)\in\mathcal{A}$ for every bank $i\in[N]$, then $\Lambda(X,m)\in\mathcal{A}$.
Proof.
Since $m\in D(X)$, we have $\Lambda(X,m)\in L^{\infty}(\mathbb{R})$ by Lemma 2.6. Suppose that $\Lambda_{i}(X,m)\in\mathcal{A}$ for every $i\in[N]$. Since the acceptance set $\mathcal{A}$ is a convex cone, we have $\sum_{i=1}^{N}\Lambda_{i}(X,m)\in\mathcal{A}$. Finally, since the underlying risk measure $\rho$ is monotone and $\sum_{i=1}^{N}\Lambda_{i}(X,m)\leq\Lambda(X,m)$, we obtain $\Lambda(X,m)\in\mathcal{A}$. ∎
Example 3.7.
Proposition 3.6 guarantees systemic acceptability by imposing a coherence assumption on the acceptance set $\mathcal{A}$. In this example, we consider a convex but not coherent risk measure to demonstrate a case where $\Lambda_{i}(X,m)\in\mathcal{A}$ for each bank $i\in[N]$ but $\Lambda(X,m)\not\in\mathcal{A}$. Specifically, we take the entropic risk measure with unit risk-aversion and the summation aggregation function over $N=2$ banks, i.e., $\mathcal{A}=\{Y\in L^{\infty}(\mathbb{R})\;|\;\mathbb{E}[1-e^{-Y}]\geq 0\}$ and $\Lambda(x,m)=\sum_{i\in[N]}(x_{i}+m_{i})$ for each $(x,m)\in\mathbb{R}^{N}\times\mathbb{R}^{N}$. Herein, we consider the natural decomposition of the aggregation function given by $\Lambda_{i}(x,m)=x_{i}+m_{i}$ for each $i\in[N]$ and $(x,m)\in\mathbb{R}^{N}\times\mathbb{R}^{N}$. Finally, we set $X=(X_{1},X_{2})^{\top}\in L^{\infty}(\mathbb{R}^{N})$ with $X_{1}=X_{2}$ a.s. and $\mathbb{P}\{X_{1}=-\frac{1}{2}\}=\mathbb{P}\{X_{1}=2\}=\frac{1}{2}$. Taking $\bar{m}=(0,0)^{\top}$, it is clear that $\mathbb{E}[1-e^{-\Lambda_{1}(X,\bar{m})}]=\mathbb{E}[1-e^{-\Lambda_{2}(X,\bar{m})}]\approx 0.108\geq 0$ whereas $\mathbb{E}[1-e^{-\Lambda(X,\bar{m})}]\approx-0.368<0$.
Corollary 3.8.
Let $(\Lambda_{i})_{i\in[N]}$ be a decomposition of $\Lambda$. Then, every $(\mathcal{A},(\Lambda_{i})_{i\in[N]})$-Nash allocation rule is acceptable.
Proof.
Let $r$ be an $(\mathcal{A},(\Lambda_{i})_{i\in[N]})$-Nash allocation rule. Let $X\in L^{\infty}_{\Lambda}(\mathbb{R}^{N})$ and $i\in[N]$. Then, there exists a decreasing sequence $(m_{i}^{n})_{n\in\mathbb{N}}$ in $\mathbb{R}$ that converges to $r_{i}(X)$ such that
$\rho(\Lambda_{i}(X,(m^{n}_{i},r_{-i}(X))))\leq 0$
for each $n\in\mathbb{N}$. By the monotonicity and upper semicontinuity of $\Lambda_{i}$, we have
[…]
Then, since $\Lambda_{i}(X,(m^{n},r_{-i}(X)))\downarrow\lim_{n\rightarrow\infty}\Lambda_{i}(X,(m^{n},r_{-i}(X)))$, we have
$\rho(\Lambda_{i}(X,r(X)))\leq\rho\left(\lim_{n\rightarrow\infty}\Lambda_{i}(X,(m^{n},r_{-i}(X)))\right)=\lim_{n\rightarrow\infty}\rho(\Lambda_{i}(X,(m^{n},r_{-i}(X))))\leq 0,$
where we use the fact that $\rho$ is continuous from above (see Assumption 3.1) and monotone. Therefore, $\Lambda_{i}(X,r(X))\in\mathcal{A}$. Following Proposition 3.6, we obtain $\Lambda(X,r(X))\in\mathcal{A}$ so that $r$ is acceptable. ∎
We wish to augment Proposition 3.6 to consider allocations when the joint distribution of the stress scenario $X$ may be unknown. Notably, in practice, estimating marginal distributions is a much simpler task than the full joint distribution, especially as the number of banks $N$ grows. While distribution estimation is not the focus of this work, the practicality of computing meaningful acceptable systemic allocations would require such considerations. The following result provides a method to find an acceptable allocation using only the marginal distributions when $\mathcal{A}$ is the acceptance set of a coherent optimized certainty equivalent risk measure; this result is adapted from Banerjee and Feinstein (2022, Lemma D.12).
Corollary 3.9.
Suppose that $\mathcal{A}$ is the acceptance set of an optimized certainty equivalent risk measure $\rho$, i.e.,
$\rho(Y):=\inf_{m\in\mathbb{R}}(m-\mathbb{E}[u(Y+m)])$
for $Y\in L^{\infty}(\mathbb{R})$, where $u\colon\mathbb{R}\to\mathbb{R}$ is defined by
$u(t)=\begin{cases}\gamma_{2}t&\text{if }t\leq 0,\\ \gamma_{1}t&\text{if }t>0,\end{cases}$
and $\gamma_{2}\geq 1\geq\gamma_{1}\geq 0$. Moreover, suppose that, for every $m\in\mathbb{R}^{N}$, the function $x\mapsto\Lambda(x,m)$ is submodular on $\mathbb{R}^{N}$, i.e.,
$\Lambda(x\vee y,m)+\Lambda(x\wedge y,m)\leq\Lambda(x,m)+\Lambda(y,m)$
for every $x,y\in\mathbb{R}^{N}$. Let $(\Lambda_{i})_{i\in[N]}$ be a decomposition of $\Lambda$. Let $X\in L^{\infty}(\mathbb{R}^{N})$ be a random shock and denote by $Z$ a comonotonic copula of $X$, i.e.,
$Z\coloneqq\left(q_{X_{1}}(U),\ldots,q_{X_{N}}(U)\right)^{\top}$
for some $U$ that is uniformly distributed on $[0,1]$, where $q_{X_{1}},\ldots,q_{X_{N}}$ are the right-continuous quantile functions of $X_{1},\ldots,X_{N}$, respectively. Let $m\in D(Z)$. If $\Lambda_{i}(Z,m)\in\mathcal{A}$ for every bank $i\in[N]$, then $\Lambda(X,m)\in\mathcal{A}$.
Proof.
The function $u$ is a utility function in the sense of Ben-Tal and Teboulle (2007, Definition 2.1) since it is concave, continuous, increasing, finitely valued with $u(0)=0$, and $1$ is a subgradient of $u$ at zero. Hence, $\rho$ is a convex risk measure by Ben-Tal and Teboulle (2007, Theorem 2.1). Moreover, the positive homogeneity of $u$ implies that $\rho$ is also coherent. Suppose that $\Lambda_{i}(Z,m)\in\mathcal{A}$ for every $i\in[N]$. First, we utilize Proposition 3.6 to conclude that $\Lambda(Z,m)\in\mathcal{A}$. We claim that $\Lambda(X,m)\in\mathcal{A}$ as well. By Shaked and Shanthikumar (2007, Theorem 9.A.21), we have $\mathbb{E}[\phi(X)]\geq\mathbb{E}[\phi(Z)]$ for every submodular function $\phi\colon\mathbb{R}^{N}\to\mathbb{R}$. In fact $x\mapsto u(\Lambda(x,m)+y)$ is submodular as it is the composition of a concave and increasing function ($u(\cdot+y)$) and a submodular and increasing function ($\Lambda(\cdot,m)$) for every $m\in D(X)$ and $y\in\mathbb{R}$. Therefore, for every $y\in\mathbb{R}$, we have $y-\mathbb{E}[u(\Lambda(X,m)+y)]\leq y-\mathbb{E}[u(\Lambda(Z,m)+y)]$. In particular, by construction of the optimized certainty equivalent, it immediately follows that $\rho(\Lambda(X,m))\leq\rho(\Lambda(Z,m))\leq 0$, i.e., $\Lambda(X,m)\in\mathcal{A}$. ∎
3.2 Existence and Uniqueness
In the above section, we take the existence of a Nash allocation rule for granted to prove that, for coherent risk measures that are continuous from above, any Nash allocation leads to an acceptable capital requirement. Within this section, we consider the existence and uniqueness of Nash allocation rules based on the self-preferential property of the decomposition of the aggregation function.
Theorem 3.10.
Let $(\Lambda_{i})_{i\in[N]}$ be a self-preferential decomposition of $\Lambda$ with constant $L\in[0,1)$.
  1. 1.
    There exists a $(\mathcal{A},(\Lambda_{i})_{i\in[N]})$-Nash allocation rule $r\colon L_{\Lambda}^{\infty}(\mathbb{R}^{N})\to\mathbb{R}^{N}$.
  2. 2.
    If $L<\frac{1}{N-1}$, then there exists a unique $(\mathcal{A},(\Lambda_{i})_{i\in[N]})$-Nash allocation rule.
Proof.
Let $X\in L_{\Lambda}^{\infty}(\mathbb{R}^{N})$ be a random shock. Recall $D(X)=D(\operatorname{ess\,inf}X)\neq\emptyset$ by Lemma 2.6(i). By self-feasibility, we can find $\bar{r}\in D(\operatorname{ess\,inf}X)$ such that
$\Lambda_{i}(\operatorname{ess\,inf}X,(\bar{r}_{i},m_{-i}))\geq 0$ (3.1)
for every $m\in D(\operatorname{ess\,inf}X)$ and $i\in[N]$. On the other hand, since $\bar{r}\in D(\operatorname{ess\,inf}X)\subseteq D(\operatorname{ess\,sup}X)$, we also have $D(\operatorname{ess\,sup}X)\neq\emptyset$. Then, by self-infeasibility, we can find $\underline{r}\in D(\operatorname{ess\,sup}X)$ such that
$\Lambda_{i}(\operatorname{ess\,sup}X,(\underline{r}_{i}-\varepsilon,m_{-i}))<0$ (3.2)
for every $m\in D(\operatorname{ess\,sup}X)$, $i\in[N]$, and $\varepsilon>0$.
For each $m\in\mathbb{R}^{N}$ and $i\in[N]$, we define
$\phi^{X}_{i}(m)\coloneqq\inf\{r_{i}\in\mathbb{R}\;|\;\Lambda_{i}(X,(r_{i},m_{-i}))\in\mathcal{A}\}.$
Let us fix $m\in D(X)=D(\operatorname{ess\,inf}X)$ and $i\in[N]$. Then, by monotonicity and (3.1),
$\phi^{X}_{i}(m)$ $\leq\inf\{r_{i}\in\mathbb{R}\;|\;\Lambda_{i}(X,(r_{i},m_{-i}))\geq 0\}$
$\leq\inf\{r_{i}\in\mathbb{R}\;|\;\Lambda_{i}(\operatorname{ess\,inf}X,(r_{i},m_{-i}))\geq 0\}\leq\bar{r}_{i}.$
Let $\varepsilon>0$. By monotonicity and (3.2),
[…]
Then, by the properties of $\rho$ as a coherent risk measure, we have
$\rho(\Lambda_{i}(X,(\underline{r}_{i}-\varepsilon,m_{-i})))\geq\rho(\Lambda_{i}(\operatorname{ess\,sup}X,(\underline{r}_{i}-\varepsilon,m_{-i})))=-\Lambda_{i}(\operatorname{ess\,sup}X,(\underline{r}_{i}-\varepsilon,m_{-i}))>0,$
Hence, $\Lambda_{i}(X,(\underline{r}_{i}-\varepsilon,m_{-i}))\notin\mathcal{A}$ and, by monotonicity, we have $\phi^{X}_{i}(m)\geq\underline{r}_{i}-\varepsilon$. Since $\varepsilon>0$ is arbitrary, we conclude that $\phi^{X}_{i}(m)\geq\underline{r}_{i}$. Therefore, $\phi^{X}(m)\in[\underline{r},\bar{r}]$ for every $m\in D(X)$. In particular, since $\bar{r}\in D(\operatorname{ess\,inf}X)=D(X)$, we have $\phi^{X}(\bar{r})\in[\underline{r},\bar{r}]$. On the other hand, since $\phi^{X}$ is decreasing by construction, we have $m\leq\bar{r}$ implies $\phi^{X}(\bar{r})\leq\phi^{X}(m)$.
We claim that $\phi^{X}(\bar{r})\in D(X)$ so that $[\phi^{X}(\bar{r}),\bar{r}]\subseteq D(X)$. The rectangular structure of $D(X)=D(\operatorname{ess\,inf}X)$ ensures that $D(X)=\bigtimes_{i=1}^{N}A_{i}$, where, for each $i\in[N]$, we have $A_{i}=\mathbb{R}$ or $A_{i}=[\bar{m}_{i},+\infty)$ for some $\bar{m}_{i}\in\mathbb{R}$. Let us fix $i\in[N]$. We may write
$\phi^{X}_{i}(\bar{r})=\inf\{r_{i}\in A_{i}\mid\Lambda_{i}(X,(r_{i},\bar{r}_{-i}))\in\mathcal{A}\}.$
Since $\phi^{X}_{i}(\bar{r})\geq\underline{r}_{i}>-\infty$ and $\rho$ is continuous from above, the above infimum is indeed a minimum so that $\phi^{X}_{i}(\bar{r})\in A_{i}$ as well. Hence, $\phi^{X}(\bar{r})\in D(\operatorname{ess\,inf}X)=D(X)$. This finishes the proof of the claim. Therefore, $\phi^{X}$ maps the compact subset $[\phi^{X}(\bar{r}),\bar{r}]$ of $D(X)$ into itself.
Next, we prove that $\phi^{X}$ is Lipschitz continuous on $D(X)$ with constant $L(N-1)$ with respect to the $\ell^{\infty}$-norm on $\mathbb{R}^{N}$. Let us fix two capital levels $m,m^{\prime}\in D(X)$ and a bank $i\in[N]$. Without loss of generality, we assume that $\phi^{X}_{i}(m)\leq\phi^{X}_{i}(m^{\prime})$. We claim that
$\phi^{X}_{i}(m^{\prime})\leq\phi^{X}_{i}(m)+L\sum_{j\in[N]\setminus\{i\}}|m_{j}-m^{\prime}_{j}|.$ (3.3)
To that end, let us fix $r_{i}\in\mathbb{R}$ such that $\Lambda_{i}(X,(r_{i},m_{-i}))\in\mathcal{A}$ and show that
$\phi^{X}_{i}(m^{\prime})\leq r_{i}+L\sum_{j\in[N]\setminus\{i\}}|m_{j}-m^{\prime}_{j}|.$
By the definition of $\phi^{X}_{i}(m^{\prime})$, this inequality follows once we establish that
$\Lambda_{i}\left(X,\left(r_{i}+L\sum_{j\in[N]\setminus\{i\}}|m_{j}-m^{\prime}_{j}|,m^{\prime}_{-i}\right)\right)\in\mathcal{A}.$ (3.4)
Let us define $m^{\prime\prime}\in\mathbb{R}^{N}$ by $m^{\prime\prime}_{j}:=m^{\prime}_{j}+|m_{j}-m^{\prime}_{j}|$ for each $j\in[N]$. Using the monotonicity and self-preferential property (for each $j\in[N]\setminus\{i\}$) of $\Lambda_{i}$, we get
$\Lambda_{i}(X,(r_{i},m_{-i}))\leq\Lambda_{i}(X,(r_{i},m^{\prime\prime}_{-i}))\leq\Lambda_{i}\left(X,\left(r_{i}+L\sum_{j\in[N]\setminus\{i\}}|m_{j}-m^{\prime}_{j}|,m^{\prime}_{-i}\right)\right).$
Since we have $\Lambda_{i}(X,(r_{i},m_{-i}))\in\mathcal{A}$ by assumption, we get (3.4) and hence (3.3). As the latter holds for every $i\in[N]$, it follows that
$\max_{i\in[N]}|\phi^{X}_{i}(m)-\phi^{X}_{i}(m^{\prime})|$ $\leq L\max_{i\in[N]}\sum_{j\in[N]\setminus\{i\}}|m_{j}-m^{\prime}_{j}|$
$\leq(N-1)L\max_{i\in[N]}\max_{j\in[N]\setminus\{i\}}|m_{j}-m^{\prime}_{j}|=(N-1)L\max_{i\in[N]}|m_{i}-m^{\prime}_{i}|$
for every $m,m^{\prime}\in D(X)$. Hence, $\phi^{X}$ is Lipschitz continuous on $D(X)$ with constant $L(N-1)$. Then, by Brouwer fixed point theorem, $\phi^{X}$ has a fixed point $r(X)\in[\phi^{X}(\bar{r}),\bar{r}]$.
Let $m\in\mathbb{R}^{N}$ be an arbitrary fixed point of $\phi^{X}$. Then, as argued for $\phi^{X}(\bar{r})$ above, the definition of $\phi^{X}(m)$ and the finiteness of $m$ guarantee that $\phi^{X}(m)\in D(X)$. Hence, $m=\phi^{X}(m)\in[\underline{r},\bar{r}]$. Then, since $\phi^{X}$ is decreasing, we have $\phi^{X}(\bar{r})\leq\phi^{X}(m)=m$. Hence, $m\in[\phi^{X}(\bar{r}),\bar{r}]$. This shows that every fixed point of $\phi^{X}$ must belong to $[\phi^{X}(\bar{r}),\bar{r}]$. With this observation, for the uniqueness argument, let us suppose that $L<\frac{1}{N-1}$. Then, $\phi^{X}$ is Lipschitz continuous on $D(X)$ with constant $L(N-1)<1$ so that it is a strict contraction. Hence, it has a unique fixed point $r(X)$ in the complete metric space $[\phi^{X}(\bar{r}),\bar{r}]\subseteq D(X)$ by Banach fixed point theorem. As every fixed point of $\phi^{X}$ belongs to this metric space, the uniqueness of a Nash allocation rule follows. ∎
Remark 3.11.
Following Theorem 3.10(2), in the setting of $N=2$ banks, there always exists a unique $(\mathcal{A},(\Lambda_{1},\Lambda_{2}))$-Nash allocation rule for a self-preferential decomposition $(\Lambda_{1},\Lambda_{2})$ of $\Lambda$. However, as the number of banks $N$ in the system grow, the uniqueness condition of Theorem 3.10(2) becomes more restrictive.
3.3 Application to Single-Element Aggregation Functions
In this section, we consider applications of Theorem 3.10 to the specific systemic risk measure structures of Example 2.9. In so doing, we propose decomposition structures that can be used for different systemic risk measure structures based on single-element aggregation functions, e.g., the ones provided in Example 2.10.
Throughout this section, we fix a single-element aggregation function $\bar{\Lambda}\colon\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$. We first introduce a refinement of decompositions in Definition 3.2 that is applicable to $\bar{\Lambda}$.
Definition 3.12.
Let $\bar{\Lambda}_{1},\dots,\bar{\Lambda}_{N}:\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ be single-element aggregation functions. The collection $(\bar{\Lambda}_{i})_{i\in[N]}$ is called a single-element decomposition of $\bar{\Lambda}$ if the following properties hold:
  1. 1.
    Consistent domain: For each $i\in[N]$, it holds $\operatorname{dom}\bar{\Lambda}_{i}=\operatorname{dom}\bar{\Lambda}$.
  2. 2.
    Dominated by $\bar{\Lambda}$: It holds $\sum_{i\in[N]}\bar{\Lambda}_{i}(x)\leq\bar{\Lambda}(x)$ for every $x\in\mathbb{R}^{N}$.
Example 3.13.
Let us revisit the single-element aggregation functions in Example 2.10 and propose single-element decompositions for them.
  1. 1.
    Sum of univariate utilities: For the single-element aggregation function $\bar{\Lambda}^{U}$ in Example 2.10(1) and Example 3.4, for each $i\in[N]$ and $x\in\mathbb{R}^{N}$, let $\bar{\Lambda}_{i}^{U}(x)\coloneqq u_{i}(x_{i})$ if $x\in\operatorname{dom}\bar{\Lambda}^{U}$ and $\bar{\Lambda}_{i}^{U}(x)\coloneqq-\infty$ otherwise. Clearly, $\bar{\Lambda}_{1}^{U},\ldots,\bar{\Lambda}_{N}^{U}$ are upper semicontinuous since $u_{1},\ldots,u_{N}$ are upper semicontinuous functions. Moreover, we have $\sum_{i\in[N]}\bar{\Lambda}^{U}_{i}(x)=\bar{\Lambda}^{U}(x)$ for every $x\in\mathbb{R}^{N}$. Hence, $(\bar{\Lambda}^{U}_{i})_{i\in[N]}$ is a single-element decomposition of $\bar{\Lambda}^{U}$.
  2. 2.
    Fund of funds: For the single-element aggregation function $\bar{\Lambda}^{FoF}$ in Example 2.10(2), for each $i\in[N]$ and $x\in\mathbb{R}^{N}$, let $\bar{\Lambda}^{FoF}_{i}(x)\coloneqq x_{i}-\gamma_{i}x_{i}^{+}-\gamma(x_{i}-\gamma_{i}x_{i}^{+})^{+}$. Then, $(\bar{\Lambda}^{FoF}_{i})_{i\in[N]}$ is a single-element decomposition of $\bar{\Lambda}^{FoF}$. We note that, in contrast to all other decompositions within this work, the fund of funds decomposition is typically strictly dominated by $\Lambda^{FoF}$ rather than guaranteeing an exact match between $\sum_{i\in[N]}\bar{\Lambda}^{FoF}_{i}$ and $\bar{\Lambda}^{FoF}$.
  3. 3.
    Mean-field utility: Consider now the single-element aggregation function $\bar{\Lambda}^{MF}$ in Example 2.10(3). For each $i\in[N]$ and $x\in\mathbb{R}^{N}$, let $\bar{\Lambda}^{MF}_{i}(x)\coloneqq u_{i}(x_{i},\frac{1}{N}\sum_{j\in[N]}x_{j})$ if $x\in\operatorname{dom}\bar{\Lambda}^{MF}$ and $\bar{\Lambda}^{MF}_{i}(x)\coloneqq-\infty$ otherwise. Then, $(\bar{\Lambda}^{MF}_{i})_{i\in[N]}$ is a single-element decomposition of $\bar{\Lambda}^{MF}$ satisfying $\sum_{i\in[N]}\bar{\Lambda}^{MF}_{i}(x)=\bar{\Lambda}^{MF}(x)$ for each $x\in\mathbb{R}^{N}$.
3.3.1 Insensitive Systemic Risk Measures
Throughout this section, we consider the insensitive systemic risk measures discussed in Example 2.9. To that end, corresponding to the single-element aggregation function $\bar{\Lambda}\colon\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$, we define the insensitive aggregation function $\Lambda^{I}\colon\mathbb{R}^{N}\times\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ by (2.3):
$\Lambda^{I}(x,m)\coloneqq\bar{\Lambda}(x)+\sum_{i\in[N]}m_{i}$
for each $x,m\in\mathbb{R}^{N}$. Recall that we have $L^{\infty}_{\Lambda^{I}}(\mathbb{R}^{N})=L^{\infty}(\operatorname{dom}\bar{\Lambda})$ in this setting. The following corollary characterizes the unique Nash allocation rule that exists for any insensitive systemic risk measure.
Corollary 3.14.
Let $(\bar{\Lambda}_{i})_{i\in[N]}$ be a single-element decomposition of $\bar{\Lambda}$. For each $i\in[N]$, define a function $\Lambda^{I}_{i}\colon\mathbb{R}^{N}\times\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ by
$\Lambda^{I}_{i}(x,m)\coloneqq\bar{\Lambda}_{i}(x)+m_{i}$
for every $x,m\in\mathbb{R}^{N}$. Then, $(\Lambda_{i}^{I})_{i\in[N]}$ is a self-preferential decomposition of $\Lambda^{I}$ with $L=0$. Moreover, the unique $(\mathcal{A},(\Lambda_{i}^{I})_{i\in[N]})$-Nash allocation rule $r^{I}\colon L^{\infty}(\operatorname{dom}\bar{\Lambda})\to\mathbb{R}^{N}$ is given by
$r^{I}(X):=\left(\rho(\bar{\Lambda}_{1}(X)),\ldots,\rho(\bar{\Lambda}_{N}(X))\right)^{\top}$
for every $X\in L^{\infty}(\operatorname{dom}\bar{\Lambda})$.
Proof.
It is easy to verify that $(\Lambda_{i}^{I})_{i\in[N]}$ is a self-preferential decomposition of $\Lambda^{I}$ with $L=0$. Then, by Theorem 3.10(2), there exists a unique Nash allocation rule $r^{I}\colon L^{\infty}(\operatorname{dom}\bar{\Lambda})\to\mathbb{R}^{N}$. It remains to show that $r^{I}$ has the form provided in the statement of the corollary. Let $X\in L^{\infty}(\operatorname{dom}\bar{\Lambda})$. Then, a $(\mathcal{A},(\Lambda_{i}^{I})_{i\in[N]})$-Nash allocation rule $r^{I}\colon L^{\infty}(\operatorname{dom}\bar{\Lambda})\to\mathbb{R}^{N}$ must satisfy
$r^{I}_{i}(X)$ $=\inf\left\{m_{i}\in\mathbb{R}\;|\;\Lambda^{I}_{i}(X,(m_{i},r^{I}_{-i}(X)))\in\mathcal{A}\right\}$
$=\inf\left\{m_{i}\in\mathbb{R}\;|\;\bar{\Lambda}_{i}(X)+m_{i}\in\mathcal{A}\right\}=\rho(\bar{\Lambda}_{i}(X)),$
which follows trivially from the primal representation of the risk measure $\rho$ (Proposition 2.2) and the definition of $\Lambda_{i}^{I}$. ∎
3.3.2 Sensitive Systemic Risk Measures
We now turn our attention to the more complex interconnections within the sensitive systemic risk measures, that is, we consider the sensitive aggregation function $\Lambda^{S}\colon\mathbb{R}^{N}\times\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ defined by (2.4):
$\Lambda^{S}(x,m)\coloneqq\bar{\Lambda}(x+m)$
for each $x,m\in\mathbb{R}^{N}$. Recall that we have $L^{\infty}_{\Lambda^{S}}(\mathbb{R}^{N})=L^{\infty}(\mathbb{R}^{N})$ in this setting.
Corollary 3.15.
Let $(\bar{\Lambda}_{i})_{i\in[N]}$ be a single-element decomposition of $\bar{\Lambda}$ that satisfies the following properties:
  1. 1.
    Self-feasible: There exists $\bar{x}\in\operatorname{dom}\bar{\Lambda}$ such that $\bar{\Lambda}_{i}(\bar{x}_{i},x_{-i})\geq 0$ for every $x\in\operatorname{dom}\bar{\Lambda}$ and $i\in[N]$.
  2. 2.
    Self-infeasible: There exists $\underline{x}\in\operatorname{dom}\bar{\Lambda}$ such that $-\infty\leq\bar{\Lambda}_{i}(\underline{x}_{i}-\varepsilon,x_{-i})<0$ for every $x\in\operatorname{dom}\bar{\Lambda}$, $i\in[N]$, and $\varepsilon>0$.
  3. 3.
    Self-preferential: There exists a constant $L\in[0,1)$ such that, for every $i,j\in[N]$ with $i\neq j$, $x\in\operatorname{dom}\bar{\Lambda}$, and $\delta\geq 0$, it holds $\bar{\Lambda}_{i}(x+L\delta e_{i})\geq\bar{\Lambda}_{i}(x+\delta e_{j})$.
For each $i\in[N]$, define a function $\Lambda^{S}_{i}\colon\mathbb{R}^{N}\times\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ by
$\Lambda^{S}_{i}(x,m)\coloneqq\bar{\Lambda}_{i}(x+m)$ (3.5)
for every $x,m\in\mathbb{R}^{N}$. Then, $(\Lambda^{S}_{i})_{i\in[N]}$ is a self-preferential decomposition of $\Lambda^{S}$. Moreover, there exists a Nash allocation rule $r^{S}\colon L^{\infty}(\mathbb{R}^{N})\to\mathbb{R}^{N}$.
Proof.
Immediately, we observe that $\Lambda^{S}_{1},\ldots,\Lambda^{S}_{N}$ are aggregation functions and $(\Lambda^{S}_{i})_{i\in[N]}$ satisfies the consistent domain property, it is dominated by $\Lambda^{S}$, and it is self-preferential. It remains to show that self-feasibility and self-infeasibility hold.
  • Self-feasible: Let $x\in\mathbb{R}^{N}$ with $D(x)\neq\emptyset$ and define $\bar{r}:=\bar{x}-x$. We have $\bar{r}\in D(x)$ since, by construction, $x+\bar{r}=\bar{x}\in\operatorname{dom}\bar{\Lambda}$. Let $m\in D(x)$, i.e., $x+m\in\operatorname{dom}\bar{\Lambda}$. Then, for every $i\in[N]$, we have
    $\Lambda_{i}^{S}(x,(\bar{r}_{i},m_{-i}))=\bar{\Lambda}_{i}(x+(\bar{r}_{i},m_{-i}))=\bar{\Lambda}_{i}(\bar{x}_{i},x_{-i}+m_{-i})\geq 0.$
  • Self-infeasible: Let $x\in\mathbb{R}^{N}$ with $D(x)\neq\emptyset$ and define $\underline{r}:=\underline{r}-x$. We have $\underline{r}\in D(x)$ since, by construction, $x+\underline{r}=\underline{x}\in\operatorname{dom}\bar{\Lambda}$. Let $m\in D(x)$, i.e., $x+m\in\operatorname{dom}\bar{\Lambda}$. Then, for every $i\in[N]$ and $\varepsilon>0$, we have
    $\Lambda_{i}^{S}(x,(\underline{r}_{i}-\varepsilon,m_{-i}))=\bar{\Lambda}_{i}(x+(\underline{r}_{i}-\varepsilon,m_{-i}))=\bar{\Lambda}_{i}(\underline{x}_{i}-\varepsilon,x_{-i}+m_{-i})<0.$
Then, by Theorem 3.10(1), the existence of a Nash allocation rule follows. ∎
As opposed to the insensitive systemic risk measures (Section 3.3.1), the Nash allocations for sensitive systemic risk measures strongly depend on the aggregation function and its chosen decomposition. To demonstrate how the Nash allocations fit within the scheme of sensitive systemic risk measures, we will revisit the single-element aggregation functions and decompositions from Example 3.13.
Example 3.16.
We consider the single-element aggregation functions and their associated decompositions in Example 3.13.
  1. 1.
    Sum of univariate utilities: Let us take $\bar{\Lambda}=\bar{\Lambda}^{U}$ and $\bar{\Lambda}_{i}=\bar{\Lambda}^{U}_{i}$ for each $i\in[N]$ as defined in Example 3.13(1). Under the assumption that $\{x_{i}\in\operatorname{dom}u_{i}\;|\;u_{i}(x_{i})\geq 0\}\not\in\{\emptyset,\mathbb{R}\}$ for each $i\in[N]$, it is easy to verify the properties in Corollary 3.15. Hence, the corresponding collection $(\Lambda^{S}_{i})_{i\in[N]}$ defined by (3.5) is a self-preferential decomposition of the sensitive aggregation function $\Lambda^{S}$, as was also argued directly in Example 3.4. As with the insensitive systemic risk measures (Section 3.3.1), there are no interactions between the capital requirements of the banks under this decomposition. Therefore, similar to the proof of Corollary 3.14, it follows that there exists a unique $(\mathcal{A},(\Lambda_{i}^{S})_{i\in[N]})$-Nash allocation rule $r^{S}\colon L^{\infty}(\mathbb{R}^{N})\to\mathbb{R}^{N}$ given by
    $r_{i}^{S}(X)=\inf\{m_{i}\in\mathbb{R}\;|\;u_{i}(X_{i}+m_{i})\in\mathcal{A}\}$
    for each $X\in L^{\infty}(\mathbb{R}^{N})$ and $i\in[N]$.
  2. 2.
    Fund of funds: Let us take $\bar{\Lambda}=\bar{\Lambda}^{FoF}$ and $\bar{\Lambda}_{i}=\bar{\Lambda}^{FoF}_{i}$ for each $i\in[N]$ as in Example 3.13(2). For this decomposition, it is easy to verify the properties of Corollary 3.15. Similar to the sum of univariate utilities, the decomposition is such that all banks model their risk independently of each other. As such, there exists a unique $(\mathcal{A},(\Lambda_{i}^{S})_{i\in[N]})$-Nash allocation rule $r^{S}\colon L^{\infty}(\mathbb{R}^{N})\to\mathbb{R}^{N}$ given by
    $r_{i}^{S}(X)=\inf\{m_{i}\in\mathbb{R}\;|\;X_{i}+m_{i}-\gamma_{i}(X_{i}+m_{i})^{+}-\gamma(X_{i}+m_{i}-\gamma_{i}(X_{i}+m_{i})^{+})^{+}\in\mathcal{A}\}$
    for each $X\in L^{\infty}(\mathbb{R}^{N})$ and $i\in[N]$.
  3. 3.
    Mean-field utility: Let us take $\bar{\Lambda}=\bar{\Lambda}^{MF}$ and $\bar{\Lambda}_{i}=\bar{\Lambda}^{MF}_{i}$ for each $i\in[N]$ as in Example 3.13(3). For simplicity, suppose that $\operatorname{dom}\bar{\Lambda}=\underline{x}+\mathbb{R}^{N}_{+}$ for some $\underline{x}\in\mathbb{R}^{N}$ and that $u_{i}$ is concave and continuously differentiable on its domain with gradient function $(\partial_{1}u_{i},\partial_{2}u_{i})$ for each $i\in[N]$. In this case, it is easy to verify that Corollary 3.15 applies if there exists some $\bar{x}\geq\underline{x}$ such that $u_{i}(\bar{x}_{i},\frac{1}{N}\bar{x}_{i}+\frac{1}{N}\sum_{j\neq i}\underline{x}_{j})\geq 0$ for every $i\in[N]$ (for self-feasibility) and
    $\inf_{\begin{subarray}{c}x\geq\underline{x}\\ \delta\geq 0\end{subarray}}\frac{\partial_{1}u_{i}\left(x_{i}+\delta,\frac{1}{N}\sum_{j\in[N]}x_{j}+\frac{1}{N}\delta\right)}{\partial_{2}u_{i}\left(x_{i},\frac{1}{N}\sum_{j\in[N]}x_{j}+\frac{1}{N}\delta\right)}>0$
    for every $i\in[N]$ (for being self-preferential). [^2] Self-infeasibility follows directly from the construction of the domain $\operatorname{dom}\bar{\Lambda}=\underline{x}+\mathbb{R}^{N}_{+}$. Because of the possible interactions between the firms in this case, it is no longer guaranteed that there exists a unique Nash allocation rule. However, despite the lack of a general theoretical guarantee, we find a unique Nash allocation for the following practical setting with $N=2$. Let us take $u_{1}=u_{2}\coloneqq\bar{u}$, where $\bar{u}(x,\bar{x})\coloneqq\log(\varepsilon+x)+\lambda\log(\varepsilon+\bar{x})$ for each $x,\bar{x}\geq 0$ and $\bar{u}(x,\bar{x})=-\infty$ otherwise, for some fixed $\varepsilon\in(0,1),\lambda>0$. Then, every $(\mathcal{A},(\Lambda_{i}^{S})_{i\in[N]})$-Nash allocation rule $r^{S}$ is uniquely evaluated at $X=0$ a.s. as $r_{1}^{S}(0)=r_{2}^{S}(0)=1-\varepsilon$.
Despite most prior examples admitting unique Nash allocation rules, it is possible to construct self-preferential decompositions with multiple Nash allocation rules. In the following example we find a continuum of Nash allocations, which highlights the importance of choosing the decomposition appropriately.
Example 3.17.
Consider a system with $N=3$ banks and a weighted summation aggregation function $\bar{\Lambda}\colon\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ defined by $\bar{\Lambda}(x)\coloneqq 1.5x_{1}+1.5x_{2}+2.25x_{3}-3$ if $x\in\mathbb{R}^{N}_{+}$ and $\bar{\Lambda}(x)\coloneqq-\infty$ otherwise. This can be considered as a special case of sum of univariate utilities decomposed as $\bar{\Lambda}^{U}_{1}(x)=u_{1}(x_{1})=1.5x_{1}-1$, $\bar{\Lambda}^{U}_{2}(x)=u_{2}(x_{2})=1.5x_{2}-1$, $\bar{\Lambda}^{U}_{3}(x)=u_{3}(x_{3})=2.25x_{3}-1$ for $x\in\mathbb{R}^{N}_{+}$ and $\bar{\Lambda}^{U}_{1}(x)=\bar{\Lambda}^{U}_{2}(x)=\bar{\Lambda}^{U}_{3}(x)=-\infty$ otherwise. This decomposition is associated with a unique associated Nash allocation rule as presented in Example 3.16(1). Let us consider instead the alternative decomposition $\bar{\Lambda}_{i}(x)=a_{i}^{\top}x-1$ for every $i\in[N]$ with $a_{1}\coloneqq(1,0,0.75)^{\top}$, $a_{2}\coloneqq(0,1,0.75)^{\top}$, $a_{3}\coloneqq(0.5,0.5,0.75)^{\top}$. In this case, for each $t\in[0,1]$, there exists a $(\mathcal{A},(\Lambda_{i}^{S})_{i\in[N]})$-Nash allocation rule $r^{S}$ such that $r^{S}(0)=(1-t,1-t,4t/3)^{\top}$. Notably, in comparison to Example 3.16(1) above, there are infinitely many Nash allocation rules with this decomposition. As a consequence, it is clear that the choice of the decomposition can greatly influence the set of Nash allocation rules in practice.
3.3.3 Scenario-Dependent Systemic Risk Measures
Finally, we consider the scenario-dependent systemic risk measures, that is, we consider the aggregation function $\Lambda^{SD}:\mathbb{R}^{N}\times\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$ defined by (2.7) for some single-element aggregation function $\bar{\Lambda}:\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$.
Proposition 3.18.
Fix a single-element aggregation function $\bar{\Lambda}:\mathbb{R}^{N}\to\mathbb{R}\cup\{-\infty\}$. If $(\Lambda_{i})_{i\in[N]}$ is a (self-preferential) decomposition of $\Lambda^{S}$, then it is also for $\Lambda^{SD}$.
Proof.
The decomposition depends on the original aggregation function only through the domination property. Since $\Lambda^{S}(x,m)\leq\Lambda^{SD}(x,m)$ for every $x,m\in\mathbb{R}^{N}$ by construction, the result holds trivially. ∎
As the decomposition constructed in Proposition 3.18 does not take advantage of scenario-dependency, we will often call $\Lambda_{i}^{SD}(x,m)\coloneqq\bar{\Lambda}_{i}(x+m)$, $x,m\in\mathbb{R}^{N}$, $i\in[N]$, the naive decomposition of $\Lambda^{SD}$. While the above observation shows that the scenario-dependent decomposition can be reduced to those found with the sensitive aggregation function, we find that constructing a scenario-dependent decomposition is non-trivial.
Remark 3.19.
We leave further considerations for the decomposition of the scenario-dependent aggregation function $\Lambda^{SD}$ for future work. For the purposes of this work, we will apply the naive decomposition so that the Nash allocation rules determined under this setting coincide exactly with those found under the sensitive systemic risk measure. For this reason, we refer the reader to Examples 3.16 and 3.17 for examples of Nash allocation rules in this systemic risk setting.
4 Case Study: Eisenberg-Noe Aggregation
For the remainder of the paper, we will consider an aggregation function derived from the Eisenberg and Noe (2001) clearing payments. In contrast to the aggregate utility functions which motivated our proposed structure, the Eisenberg-Noe aggregation function has complex dependencies between the banks within its natural decomposition. As far as the authors are aware, the Eisenberg-Noe aggregation function was first proposed in Feinstein et al. (2017), though it is a special case of the optimization aggregation function in Chen et al. (2013, Example 4) and is closely related to the contagion aggregation function in Chen et al. (2013, Example 7).
4.1 Eisenberg-Noe Aggregation and Decomposition
Consider a system of $N$ banks that are interconnected via payment obligations. For each $i,j\in[N]$, we let $\bar{p}_{ij}\geq 0$ denote the obligations that bank $i$ owes to bank $j$; we set $\bar{p}_{ii}=0$ to avoid self-dealing. In addition, each bank $i$ also has an obligation $\bar{p}_{i0}>0$ external to the financial system; we often refer to these obligations as those due to “society”. The  Eisenberg and Noe (2001) clearing system assumes that these obligations are repaid proportionally. Therefore, from this construction, we set $\pi_{ij}=\bar{p}_{ij}/\sum_{k=0}^{N}\bar{p}_{ik}$ to be the relative liability of bank $i\in[N]$ to entity (bank or society) $j\in[N]\cup\{0\}$. Assuming that the $N$ banks hold external assets worth $x\in\mathbb{R}^{N}_{+}$ and following the priority of debt over equity and limited liabilities, a clearing payment vector $p(x)\in\mathbb{R}^{N}$ is defined a solution of the fixed point equation
$p_{i}(x)=\min\left\{\sum_{j=0}^{N}\bar{p}_{ij}\;,\;x_{i}+\sum_{j=1}^{N}\pi_{ji}p_{j}(x)\right\},\quad i\in[N].$ (4.1)
Immediately, by applying Tarski fixed point theorem, there exists greatest and least clearing payment vectors for $x\in\mathbb{R}^{N}_{+}$. Moreover, as provided by Eisenberg and Noe (2001, Theorem 2), we recover a unique clearing payment vector $p(x)$ for every $x\in\mathbb{R}^{N}_{+}$ due to the strictly positive obligations to society. In the following proposition, we provide properties of the clearing payment vector.
Proposition 4.1.
Consider a financial network of $N$ banks with obligations $\bar{p}\in\mathbb{R}^{N\times(N+1)}_{+}$ satisfying $\bar{p}_{i0}>0$ for each $i\in[N]$. For every asset level $x\in\mathbb{R}^{N}_{+}$, there exists a unique clearing payment vector $p(x)\in\mathbb{R}^{N}_{+}$ satisfying (4.1). Furthermore, the function $p\colon\mathbb{R}^{N}_{+}\to\mathbb{R}^{N}_{+}$ is non-expansive [^3] That is, $\sum_{i\in[N]}|p_{i}(x)-p_{i}(x^{\prime})|\leq\sum_{i\in[N]}|x_{i}-x^{\prime}_{i}|$ for every $x,x^{\prime}\in\mathbb{R}^{N}_{+}$., concave, increasing, submodular, and it holds $p_{j}(x+\delta e_{i})\leq p_{j}(x+\delta\pi_{i\cdot})$ for every $i,j\in[N]$ with $i\neq j$, $x\in\mathbb{R}^{N}_{+}$, and $\delta\geq 0$.
Proof.
Due to the strictly positive obligations to society, the provided system is a regular network as per Eisenberg and Noe (2001, Definition 5). Therefore, by Eisenberg and Noe (2001, Theorem 2), the uniqueness of the clearing payment vector $p(x)$ follows for every $x\in\mathbb{R}^{N}_{+}$. Furthermore, by Eisenberg and Noe (2001, Lemma 5), the clearing payment function $p\colon\mathbb{R}^{N}_{+}\to\mathbb{R}^{N}_{+}$ is concave, increasing, and non-expansive; its submodularity follows from Banerjee and Feinstein (2022, Proposition A.7).
Finally, let us fix $i\in[N]$, $x\in\mathbb{R}^{N}_{+}$, and $\delta\geq 0$. Note that $p_{i}(x+\delta\pi_{i\cdot})+\delta\geq p_{i}(x)+\delta\geq p_{i}(x+\delta e_{i})$ by a simple application of the monotonicity and non-expansiveness of the clearing payment function. Let $J\coloneqq\{j\in[N]\setminus\{i\}\;|\;p_{j}(x+\delta\pi_{i\cdot})<\sum_{k=0}^{N}\bar{p}_{jk}\}$ be the collection of banks (not $i$) that fail to make their payments in full under asset level $x+\delta\pi_{i\cdot}\in\mathbb{R}^{N}_{+}$. For every $j\in J^{c}:=([N]\setminus\{i\})\setminus J$, we have $p_{j}(x+\delta\pi_{i\cdot})=\sum_{k=0}^{N}\bar{p}_{jk}\geq p_{j}(x+\delta e_{i})$ by construction. Assume that $J\neq\emptyset$ as otherwise the proof is completed. By the construction of clearing payments, with an abuse of notation for sub-vectors and sub-matrices, we have
$p_{J}(x+\delta\pi_{i\cdot})$ $=x_{J}+\pi_{iJ}^{\top}[p_{i}(x+\delta\pi_{i\cdot})+\delta]+\pi_{J^{c}J}^{\top}\sum_{k=0}^{N}\bar{p}_{J^{c}k}+\pi_{JJ}^{\top}p_{J}(x+\delta\pi_{i\cdot})$
$=(I_{|J|}-\pi_{JJ}^{\top})^{-1}\left(x_{J}+\pi_{iJ}^{\top}[p_{i}(x+\delta\pi_{i\cdot})+\delta]+\pi_{J^{c}J}^{\top}\sum_{k=0}^{N}\bar{p}_{J^{c}k}\right)$
$\geq(I_{|J|}-\pi_{JJ}^{\top})^{-1}\left(x_{J}+\pi_{iJ}^{\top}p_{i}(x+\delta e_{i})+\pi_{J^{c}J}^{\top}p_{J^{c}}(x+\delta e_{i})\right)\geq p_{J}(x).$
In this calculation, the second line follows from rearranging terms and noting, due to the obligations to society, that $I_{|J|}-\pi_{JJ}^{\top}$ is invertible. The first inequality takes advantage of the positivity of the Leontief inverse as well as bounds on $p_{k}(x+\delta\pi_{i\cdot})$ for $k\in J^{c}\cup\{i\}$. The final inequality follows directly from the construction of the clearing payments. ∎
As detailed in, e.g., Feinstein et al. (2017), the payments to society form a meaningful aggregation function for measuring systemic risk. The idea is that, a priori, a regulator does not care if a bank fails; instead, the risks are whether these failures spread to society, i.e., the real economy. Assuming that the regulator requires at least $\gamma\in(0,1)$ fraction of all obligations to society to be repaid, the corresponding Eisenberg-Noe aggregation function $\bar{\Lambda}^{EN}:\mathbb{R}^{N}\to\mathbb{R}$ is defined by
$\bar{\Lambda}^{EN}(x):=\sum_{i\in[N]}\left(\pi_{i0}p_{i}(x)-\gamma\bar{p}_{i0}\right)$
for every $x\in\mathbb{R}^{N}_{+}$, and by $\bar{\Lambda}^{EN}(x):=-\infty$ otherwise.
Immediately upon the introduction of the single-element aggregation function $\bar{\Lambda}^{EN}$, there is a natural decomposition. Specifically, rather than aggregating all payments to society together, the regulator can impose the condition that each bank $i\in[N]$ needs to cover $\gamma$ fraction of its own obligations to society, i.e., we may define
$\bar{\Lambda}^{EN}_{i}(x):=\pi_{i0}p_{i}(x)-\gamma\bar{p}_{i0}$
whenever $x\in\mathbb{R}^{N}_{+}$, and $\bar{\Lambda}^{EN}_{i}(x):=-\infty$ otherwise. It can trivially be verified that $(\bar{\Lambda}^{EN}_{i})_{i\in[N]}$ is a single-element decomposition of $\bar{\Lambda}^{EN}$ in the sense of Definition 3.12.
Proposition 4.2.
Consider a financial network of $N$ banks with obligations $\bar{p}\in\mathbb{R}^{N\times(N+1)}_{+}$ satisfying $\bar{p}_{i0}>0$ for each $i\in[N]$. The collection $(\bar{\Lambda}_{i}^{EN})_{i\in[N]}$ is a single-element decomposition of the Eisenberg-Noe aggregation function $\bar{\Lambda}^{EN}$ satisfying all properties in Corollary 3.15 with a self-preferential constant $L$ that satisfies
$L\leq\max_{\lambda\in\{0,1\}^{N}}\max_{\begin{subarray}{c}i,j\in[N]\colon\\ i\neq j\end{subarray}}\frac{[(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}]_{ij}}{[(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}]_{jj}}\leq\max_{i\in[N]}(1-\pi_{i0})<1.$
Proof.
First, we note that $\operatorname{dom}\bar{\Lambda}_{i}^{EN}=\operatorname{dom}\bar{\Lambda}^{EN}=\mathbb{R}^{N}_{+}$ for all banks $i\in[N]$, hence the consistent domain property holds. Additionally, we have $\sum_{i\in[N]}\bar{\Lambda}^{EN}_{i}=\bar{\Lambda}^{EN}$ so that $(\bar{\Lambda}^{EN}_{i})_{i\in[N]}$ is dominated by $\bar{\Lambda}^{EN}$. It remains to show that the additional properties in Corollary 3.15 hold. Self-feasibility and self-infeasibility follow with $\bar{x}_{i}:=\sum_{j=0}^{N}\bar{p}_{ij}$ and $\underline{x}_{i}:=0$ for every bank $i\in[N]$ as we have, trivially, $p_{i}(\bar{x}_{i},x_{-i})=\sum_{j=0}^{N}\bar{p}_{ij}$ and $(\underline{x}_{i}-\varepsilon,x_{-i})\not\in\mathbb{R}^{N}_{+}$ for every $x\in\mathbb{R}^{N}_{+}$ and $\varepsilon>0$. Finally, the exact formulation of the self-preferential constant $L$ comes from the worst case using the derivatives of the clearing payments as found in, e.g., Liu and Staum (2010), i.e.,
$\frac{\partial}{\partial x_{j}}p_{i}(x)=[(I_{N}-\operatorname{diag}(\mathbbm{1}_{\{p(x)<\sum_{k=0}^{N}\bar{p}_{\cdot k}\}})\pi^{\top})^{-1}]_{ij}\mathbbm{1}_{\{p_{j}(x)<\sum_{k=0}^{N}\bar{p}_{jk}\}}.$
Therefore, locally in the default region $\lambda=\mathbbm{1}_{\{p(x)<\sum_{k=0}^{N}\bar{p}_{\cdot k}\}}$,
$L(x)=\max_{\begin{subarray}{c}i\neq j\\ \lambda_{i}=1\end{subarray}}\frac{[(I_{N}-\operatorname{diag}(\lambda)\pi^{\top})^{-1}]_{ij}\lambda_{j}}{[(I_{N}-\operatorname{diag}(\lambda)\pi^{\top})^{-1}]_{ii}}=\max_{\begin{subarray}{c}i\neq j\\ \lambda_{j}=1\end{subarray}}\frac{[(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}]_{ji}\lambda_{i}}{[(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}]_{jj}}$
where the equality follows from the use of the Leontief inverses $(I_{N}-\operatorname{diag}(\lambda)\pi^{\top})^{-1}$ and $(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}$. Immediately, the first inequality becomes clear:
$L\leq\max_{\lambda\in\{0,1\}^{N}}\max_{\begin{subarray}{c}i,j\in[N]\colon\\ i\neq j\end{subarray}}\frac{[(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}]_{ij}}{[(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}]_{jj}}.$
Next, we bound this constant by $\max_{i\in[N]}(1-\pi_{i0})$.
Let us fix a binary vector $\lambda\in\{0,1\}^{N}$ and define $E(\lambda):=\{i\in[N]\;|\;\lambda_{i}=0\}$ and $E(\lambda)^{c}:=[N]\setminus E(\lambda)$. Note that $\operatorname{diag}(\lambda)\pi$ is the matrix obtained by replacing the $i^{\text{th}}$ row of $\pi$ with a row of zeroes for each $i\in E(\lambda)$. Let $\tilde{\pi}(\lambda)=(\pi_{ij}(\lambda))_{i,j\in[N]\cup\{0\}}\in\mathbb{R}^{(N+1)\times(N+1)}_{+}$ be the relative liabilities matrix of the extended network in which node 00 represents society and nodes $1,\ldots,N$ represent the banks. Since society does not have any obligations to the banks, we take $\pi_{00}(\lambda):=1$ and $\pi_{0j}(\lambda):=0$ for each $j\in[N]$. Moreover, in this network, we assume that all banks in $E(\lambda)$ have their full obligations to society, i.e.,
$\pi_{ij}(\lambda)=\begin{cases}1&\text{if }i\in E(\lambda)\cup\{0\},\ j=0,\\ 0&\text{if }i\in E(\lambda)\cup\{0\},\ j\in[N],\\ \pi_{ij}&\text{if }i\in E(\lambda)^{c},\ j\in[N]\cup\{0\}.\end{cases}$
Note that $\tilde{\pi}(\lambda)$ is a right stochastic matrix and we may view it as the one-step transition probability matrix of a time-homogeneous Markov chain with state space $[N]$. Clearly, state 00 is the only recurrent state and all other states are transient. Moreover, since state 00 is absorbing, given $i,j\in[N]$, there is no path from $i$ to $j$ that goes through state 00; in particular, the corresponding multi-step transition probabilities can be calculated by only using the original matrix $\operatorname{diag}(\lambda)\pi$ via
$[(\operatorname{diag}(\lambda)\pi)^{n}]_{ij}=[(\tilde{\pi}(\lambda))^{n}]_{ij},\quad n\in\mathbb{Z}_{+}.$
Moreover, we have
$(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}=\sum_{n=0}^{\infty}(\operatorname{diag}(\lambda)\pi)^{n},\quad(I_{N+1}-\tilde{\pi}(\lambda))^{-1}=\sum_{n=0}^{\infty}(\tilde{\pi}(\lambda))^{n}.$
Hence, for every $i,j\in[N]$, we also have
$[(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}]_{ij}=[(I_{N+1}-(\tilde{\pi}(\lambda))^{-1}]_{ij}.$
The matrix $a(\lambda)=(a_{ij}(\lambda))_{i,j\in[N]\cup\{0\}}:=(I_{N+1}-\tilde{\pi}(\lambda))^{-1}$ has a well-known interpretation in the theory of Markov chains. For each $j\in[N]\cup\{0\}$, let $N_{j}(\lambda)$ be the total number of visits of the chain to state $j$. Then, for each $i\in[N]\cup\{0\}$, we have $a_{ij}(\lambda)=\mathbb{E}_{i}[N_{j}(\lambda)]$, where $\mathbb{E}_{i}$ denotes the conditional expectation given that the chain starts at state $i$ at time zero. Since state 00 is the unique absorbing state, we have $a_{0j}(\lambda)=0$ for every $j\in[N]$ and $a_{i0}(\lambda)=+\infty$ for every $i\in[N]\cup\{0\}$.
For every $i,j\in[N]$ with $i\neq j$, by Çınlar (1975, Corollary 5.2.13), we have
$a_{ij}(\lambda)=\mathbb{P}_{i}\{T_{j}(\lambda)<+\infty\}a_{jj}(\lambda),$
where $T_{j}(\lambda)$ is the first time that the chain visits state $j$ after time zero and $\mathbb{P}_{i}$ denotes the conditional probability given that the chain starts at state $i$ at time zero. Note that
$\mathbb{P}_{i}\{T_{j}(\lambda)=+\infty\}\geq\pi_{i0}(\lambda)=\begin{cases}1&\text{if }i\in E(\lambda),\\ \pi_{i0}&\text{if }i\in E(\lambda)^{c}.\end{cases}$
since state 00 is absorbing. Hence, $\mathbb{P}_{i}\{T_{j}(\lambda)<+\infty\}=0$ if $i\in E(\lambda)$ and $\mathbb{P}_{i}\{T_{j}(\lambda)<+\infty\}\leq 1-\pi_{i0}$ if $i\in E(\lambda)^{c}$. Then, we get
$\max_{\begin{subarray}{c}i,j\in[N]\colon\\ i\neq j\end{subarray}}\frac{[(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}]_{ij}}{[(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}]_{jj}}$ $=\max_{\begin{subarray}{c}i,j\in[N]\colon\\ i\neq j\end{subarray}}\frac{[(I_{N+1}-\tilde{\pi}(\lambda))^{-1}]_{ij}}{[(I_{N+1}-\tilde{\pi}(\lambda))^{-1}]_{jj}}=\max_{\begin{subarray}{c}i,j\in[N]\colon\\ i\neq j\end{subarray}}\frac{a_{ij}(\lambda)}{a_{jj}(\lambda)}$
$\leq\max_{\begin{subarray}{c}i,j\in[N]\colon\\ i\neq j\end{subarray}}\mathbb{P}_{i}\{T_{j}(\lambda)<+\infty\}\leq\max_{i\in E(\lambda)^{c}}(1-\pi_{i0}).$
Therefore,
$L\leq\max_{\lambda\in\{0,1\}^{N}}\max_{\begin{subarray}{c}i,j\in[N]\colon\\ i\neq j\end{subarray}}\frac{[(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}]_{ij}}{[(I_{N}-\operatorname{diag}(\lambda)\pi)^{-1}]_{jj}}\leq\max_{\lambda\in\{0,1\}^{N}}\max_{i\in E(\lambda)^{c}}(1-\pi_{i0})=\max_{i\in E\setminus\{0\}}(1-\pi_{i0})<1$
since we assume that $\pi_{i0}>0$ for every $i\in[N]$. ∎
4.2 Nash Allocation
Given the Eisenberg-Noe aggregation function $\bar{\Lambda}^{EN}$ and its natural decomposition $(\bar{\Lambda}_{i}^{EN})_{i\in[N]}$, we wish to consider various properties of the associated Nash allocations for sensitive systemic risk measures. We note that the Nash allocations for the scenario-dependent systemic risk measures coincide exactly with those of the sensitive systemic risk measures (Proposition 3.18) whereas the Nash allocations for the insensitive systemic risk measures can trivially be found (Corollary 3.14).
As we only consider the sensitive systemic risk measures within this section, we define
$\Lambda^{EN}(x,m)=\bar{\Lambda}^{EN}(x+m),\quad\Lambda_{i}^{EN}(x,m)=\bar{\Lambda}_{i}^{EN}(x+m)$
for each $x,m\in\mathbb{R}^{N}$ and $i\in[N]$ throughout. We also fix an acceptance set $\mathcal{A}$ satisfying Assumption 3.1. First, within Corollary 4.3, we summarize the existence of a Nash allocation. We then consider a single optimization problem which can be used to find a Nash allocation for the natural decomposition. Finally, we conclude with a quick discussion on two situations in which the Nash allocation is unique.
Corollary 4.3.
Consider a financial network of $N$ banks with obligations $\bar{p}\in\mathbb{R}^{N\times(N+1)}_{+}$ satisfying $\bar{p}_{i0}>0$ for each $i\in[N]$.
  1. 1.
    There exists a $(\mathcal{A},(\Lambda_{i}^{EN})_{i\in[N]})$-Nash allocation rule $r^{*}\colon L^{\infty}(\mathbb{R}^{N})\to\mathbb{R}^{N}$.
  2. 2.
    Suppose that $\mathcal{A}$ is the acceptance set of a coherent optimized certainty equivalent risk measure as in Corollary 3.9. Let $X\in L^{\infty}(\mathbb{R}^{N})$ be a stress scenario and denote by $Z^{X}$ the comonotonic copula of $X$. Then, $r^{*}(Z^{X})\in\mathbb{R}^{N}$ is acceptable under $X$, i.e., $\bar{\Lambda}^{EN}(X+r^{*}(Z^{X}))\in\mathcal{A}$.
Proof.
The existence of a $(\mathcal{A},(\Lambda_{i}^{EN})_{i\in[N]})$-Nash allocation rule follows from Corollary 3.15 as the natural decomposition satisfies all properties of that result (Proposition 4.2). When $\mathcal{A}$ is the acceptance set of a coherent optimized certainty equivalent risk measure as in Corollary 3.9, by the submodularity of the Eisenberg-Noe aggregation function (Proposition 4.1), the acceptability of the comonotonic Nash allocation $r^{*}(Z^{X})$ follows from Corollary 3.9. ∎
Though we only guarantee the existence of a Nash allocation in Corollary 4.3, we are able to provide a constructive approach to compute one such allocation. In particular, the computed allocation will be the one that has the minimal total capital allocation. To construct this allocation, we are inspired by Braouezec and Kiani (2023, Proposition 2); in contrast to that result, we are not considering the shared constraint game; however, herein, the same approach can be used to find a $(\mathcal{A},(\Lambda_{i}^{EN})_{i\in[N]})$-Nash allocation with its individual constraints. Notably, under a finite probability space with coherent (polyhedral) acceptance set, the approach provided in Theorem 4.4 below only requires solving a single convex (resp. linear) program.
Theorem 4.4.
Consider a financial network of $N$ banks with obligations $\bar{p}\in\mathbb{R}^{N\times(N+1)}_{+}$ satisfying $\bar{p}_{i0}>0$ for each $i\in[N]$. For each $X\in L^{\infty}(\mathbb{R}^{N})$, there exists a solution to the optimization problem
$R^{*}(X):=\operatorname*{arg\,min}_{m\in\mathbb{R}^{N}}\left\{\sum_{i\in[N]}m_{i}\;|\;\bar{\Lambda}_{i}^{EN}(X+m)\in\mathcal{A}\,\forall i\in[N]\right\}\neq\emptyset.$ (4.2)
Moreover, let $r^{*}:L^{\infty}(\mathbb{R}^{N})\to\mathbb{R}^{N}$ be a selector of $R^{*}$ (i.e., $r^{*}(X)\in R^{*}(X)$ for every $X\in L^{\infty}(\mathbb{R}^{N})$), then $r^{*}$ is a $(\mathcal{A},(\Lambda_{i}^{EN})_{i\in[N]})$-Nash allocation rule.
Proof.
Let $X\in L^{\infty}(\mathbb{R}^{N})$. We first argue that $R^{*}(X)\neq\emptyset$. By Theorem 3.10, there exists a Nash allocation rule $r\colon L^{\infty}(\mathbb{R}^{N})\to\mathbb{R}^{N}$. In particular, $\bar{\Lambda}^{EN}_{i}(X+r(X))\in\mathcal{A}$ for each $i\in[N]$. Consequently, the feasible region of the problem in (4.2) is a non-empty set. Moreover, it is also a closed set, which can be checked similar to the proof of Corollary 3.8. Next, for a feasible solution $m\in\mathbb{R}^{N}$, the acceptance set constraints implicitly imply that $X+m\in\operatorname{dom}\bar{\Lambda}^{EN}=\mathbb{R}^{N}_{+}$, i.e., $m_{i}\geq-\operatorname{ess\,inf}X_{i}$ for every $i\in[N]$. Let $v(X)$ denote the optimal value of the problem. Then, we have $v(X)\in\mathbb{R}$ and
$v(X)=\negthinspace\inf_{m\in\mathbb{R}^{N}}\negthinspace\left\{\sum_{i\in[N]}m_{i}\;|\;\bar{\Lambda}^{EN}_{i}(X+m)\in\mathcal{A}\ \forall i\in[N],-\operatorname{ess\,inf}X\leq m,\sum_{i\in[N]}m_{i}\leq\sum_{i\in[N]}\bar{m}_{i}(X)\right\},$
where $\bar{m}(X)\in\mathbb{R}^{N}$ is an arbitrarily fixed feasible solution of the original problem. Since the feasible region of the new problem is compact and the objective function is continuous, there exists an optimal solution $r^{*}(X)$, which is also an optimal solution for the original problem.
Next, we show that the functional $r^{*}\colon L^{\infty}(\mathbb{R}^{N})\to\mathbb{R}^{N}$ defined above is a Nash allocation rule. Note that a vector $m\in\mathbb{R}^{N}$ satisfies the definition of a Nash allocation rule at $X\in L^{\infty}(\mathbb{R}^{N})$ if and only if, for every $i\in[N]$, we have $\rho(\bar{\Lambda}^{EN}_{i}(X+m))\leq 0$ and the following implication holds:
$\rho(\bar{\Lambda}^{EN}_{i}(X+m))<0\quad\Rightarrow\quad m_{i}=-\operatorname{ess\,inf}X_{i}.$
To get a contradiction, suppose that $r^{*}(X)$ does not satisfy this definition. Then, using the notation of $\phi^{X}$ from the proof of Theorem 3.10, there exists some $i\in[N]$ such that $m^{*}_{i}(X):=\phi^{X}_{i}(r^{*}(X))<r_{i}^{*}(X)$. Let $\delta\in(0,r_{i}^{*}(X)-m^{*}_{i}(X)]$. Therefore, for every $j\in[N]\setminus\{i\}$, we have
$\rho(\bar{\Lambda}_{j}^{EN}(X+r^{*}(X)+\delta\pi_{i\cdot}-\delta e_{i}))\leq\rho(\bar{\Lambda}_{j}^{EN}(X+r^{*}(X)))\leq 0$
by the final property in Proposition 4.1. Additionally, by the monotonicity of the Eisenberg-Noe decomposition, the choice of $\delta$, and the definition of $m^{*}_{i}(X)$, we have
$\rho(\bar{\Lambda}_{i}^{EN}(X+r^{*}(X)+\delta(\pi_{i\cdot}-e_{i})))\leq\rho(\bar{\Lambda}_{i}^{EN}(X+r^{*}(X)-\delta e_{i}))\leq\rho(\bar{\Lambda}_{i}^{EN}(X+(\bar{m}^{*}_{i}(X),r_{-i}^{*}(X)))\leq 0.$
Therefore, $r^{*}(X)+\delta(\pi_{i\cdot}-e_{i})$ is feasible for the problem in (4.2) and we have $\sum_{j\in[N]}(r_{j}^{*}(X)+\delta(\pi_{ij}-e_{ij}))=\sum_{j\in[N]}r_{j}^{*}(X)-\pi_{i0}<\sum_{j\in[N]}r_{j}^{*}(X)$ since $\pi_{i0}>0$, contradicting the optimality of $r^{*}(X)$. ∎
Thus far, we have considered the existence of Nash allocation rules (Corollary 4.3) and formulated an optimization problem to compute one such rule (Theorem 4.4). We conclude our general discussion on the Nash allocation rules for the sensitive Eisenberg-Noe systemic risk measure by providing two cases in which uniqueness of the Nash allocation can be guaranteed.
Corollary 4.5.
Consider a financial network of $N$ banks with obligations $\bar{p}\in\mathbb{R}^{N\times(N+1)}_{+}$ satisfying $\bar{p}_{i0}>0$ for each $i\in[N]$.
  1. 1.
    If $\pi_{i0}>\frac{N-2}{N-1}$ for every bank $i\in[N]$, then there exists a unique $(\mathcal{A},(\Lambda_{i}^{EN})_{i\in[N]})$-Nash allocation rule.
  2. 2.
    Consider a deterministic shock $X=x$, where $x\in\mathbb{R}^{N}$. Then, every $(\mathcal{A},(\Lambda_{i}^{EN})_{i\in[N]})$-Nash allocation rule takes the same value at $X$.
Proof.
The first claim is a trivial consequence of Theorem 3.10(2) using the bound on the self-preferential constant $L\leq\max_{i\in[N]}(1-\pi_{i0})$ from Proposition 4.2.
To prove the second statement, let $X=x$ with $x\in\mathbb{R}^{N}$. Let $i\in[N]$. By construction, $\rho(\bar{\Lambda}_{i}^{EN}(x+m))=-\bar{\Lambda}_{i}^{EN}(x+m)$ for every $m\geq-x$. Therefore, the acceptability constraint for bank $i$ is equivalent to $p_{i}(x+m)\geq\gamma\sum_{j=0}^{N}\bar{p}_{ij}$. We take advantage of this structure in order to provide the following algorithm to compute a Nash allocation $m(x)\in\mathbb{R}^{N}$:
  1. 1.
    Initialize $k\leftarrow 0$ and set $I^{0}\coloneqq\emptyset$.
  2. 2.
    Define $p_{i}^{k}\coloneqq\gamma\sum_{j=0}^{N}\bar{p}_{ij}$ if $i\not\in I^{k}$. Define $(p_{i}^{k})_{i\in I^{k}}$ as the unique solution of the Eisenberg-Noe clearing problem
    $p_{i}^{k}=\min\left\{\sum_{j=0}^{N}\bar{p}_{ij}\;,\;\gamma\sum_{j\not\in I^{k}}\bar{p}_{ji}+\sum_{j\in I^{k}}\pi_{ji}p_{j}^{k}\right\}\quad\forall i\in I^{k}.$
  3. 3.
    Define $m^{k}$ by
    $m_{i}^{k}\coloneqq\begin{cases}-x_{i}+\gamma\sum_{j=0}^{N}\bar{p}_{ij}-\sum_{j=1}^{N}\pi_{ji}p_{j}^{k}&\text{if }i\not\in I^{k}\\ -x_{i}&\text{if }i\in I^{k}.\end{cases}$
  4. 4.
    Define $I^{k+1}\coloneqq\{i\in[N]\;|\;m_{i}^{k}\leq-x_{i}\}$ and iterate $k\leftarrow k+1$.
  5. 5.
    If $I^{k}=I^{k-1}$, then terminate the algorithm with $m(x)\coloneqq m^{k-1}$ and $I^{*}\coloneqq I^{k-1}$. Otherwise, return to step (2).
Note that this algorithm converges in at most $N$ iterations as $I^{k}\supseteq I^{k-1}$ for every $k\geq 1$. Furthermore, by construction, it is clear that $m(x)$ is a Nash allocation, i.e., $m_{i}(x)=\inf\{m_{i}\in\mathbb{R}\mid\bar{\Lambda}^{EN}_{i}(x+(m_{i},m_{-i}(x)))\in\mathcal{A}\}$ for every $i\in[N]$, as no bank can reduce their capital requirement without violating the acceptability constraint.
Finally, let $r\colon L^{\infty}(\mathbb{R}^{N})\to\mathbb{R}^{N}$ be a Nash allocation rule. As in the above algorithm, $r(x)$ can be associated with the set of banks with no capital requirement $I:=\{i\in[N]\;|\;r_{i}(x)\leq-x_{i}\}$. By the construction of $I^{*}$ to be the smallest possible set of banks who can passively maintain acceptability, it follows that $I\supseteq I^{*}$. Assume $I\supsetneq I^{*}$ and fix some $i\in I\backslash I^{*}$, i.e., $r_{i}(x)=-x_{i}<m_{i}(x)$. However, by simple inspection, $p_{i}(x+r)<\gamma\sum_{j=0}^{N}\bar{p}_{ij}$ (or else $i\in I^{*}$ must hold). Therefore $I=I^{*}$ and, as a consequence, $r(x)=m(x)$, completing the proof. ∎
Remark 4.6.
Though Corollary 4.5 provides sufficient conditions for the uniqueness of a Nash allocation, the authors have not been able to construct any examples with multiple Nash allocations for the Eisenberg-Noe aggregation function. This hints that a stronger result may be possible, but we leave that for future work.
4.3 Numerical Examples
We conclude our discussion of the Eisenberg-Noe aggregation function by presenting two detailed case studies where we compute explicitly the Nash allocations under given stress scenarios. In both examples, we choose settings from Corollary 4.5 that guarantee the uniqueness of the Nash allocations: first, in Section 4.3.1, we consider the $N=2$-bank setting; second, in Section 4.3.2, we consider a deterministic setting calibrated to data from the European Banking Authority (EBA). Throughout, we compare the Nash allocations with the minimal possible systemically acceptable capital requirements. In all cases, though the Nash allocations may require strictly larger total capital, the Nash allocations provide financially justifiable selections from the set of all acceptable capitals.
4.3.1 Two-Bank Example
Consider the $N=2$-bank system with obligations $\bar{p}_{10}=\bar{p}_{20}=\bar{p}_{12}=1$, $\bar{p}_{21}=\frac{1}{2}$, and $\bar{p}_{11}=\bar{p}_{22}=0$. Furthermore, we set acceptability such that $\gamma=95\%$ of obligations to society are recovered. Recall from Corollary 4.5 that this network immediately admits a unique Nash allocation for any stress scenario and coherent risk measure. For the numerics, we set $X_{i}\sim U_{i}\sum_{j=0}^{N}\bar{p}_{ij}$ for each bank $i\in[N]$, where $U=(U_{1},\ldots,U_{N})^{\top}$ is a vector of standard uniform random variables generated from a Gaussian copula with correlations $\frac{1}{2}$. [^4] The stress scenario is implemented with $1000$ sample values. Finally, we will consider two risk measures within this example:
  1. 1.
    Expectation risk measure: $\mathcal{A}^{\mathbb{E}}:=\{Y\in L^{\infty}(\mathbb{R})\;|\;\mathbb{E}[Y]\geq 0\}$; and
  2. 2.
[…]
Insensitive Systemic Risk Measure
Expectation Risk Measure Average Value-at-Risk
Minimal Nash Comonotonic Minimal Nash Comonotonic
Bank 1 0.2772 0.2865 0.8723 0.9088
Bank 2 0.1657 0.1996 0.8299 0.8946
Total 0.4429 0.4429 0.4861 1.6867 1.7022 1.8034
Sensitive Systemic Risk Measure
Expectation Risk Measure Average Value-at-Risk
Minimal Nash Comonotonic Minimal Nash Comonotonic
Bank 1 0.8994 0.9201 0.9403 1.3311 1.3569 1.3703
Bank 2 0.1697 0.1502 0.2192 0.3963 0.4024 0.4331
Total 1.0691 1.0703 1.1595 1.7275 1.7593 1.8034
SD 0.9787 1.6402
Table 1: Section 4.3.1: Summary table of capital requirements under different systemic risk measures and utilizing (i) minimal capital requirements; (ii) the Nash allocation rule; and (iii) the Nash allocation rule under the comonotonic copula.
Within Table 1, we can see how the choice of aggregation and acceptance set alter the capital requirements assessed. Throughout, we take the minimal capital requirements as the minimizer of $\{m_{1}+m_{2}\;|\;m\in R(X)\}$ for the appropriate choice of systemic risk measure $R$. Notably, for the insensitive and scenario-dependent (SD) cases, because these systemic risk measures are (truncated) half-spaces, we do not record the individual minimal requirements but only the total capital requirement assessed. In fact, as we take the naive decomposition for the scenario-dependent setting, we do not display the specific Nash allocations explicitly, but rather consider it to be a sub-case of the sensitive systemic risk measures. In addition, we benchmark these results by displaying the Nash allocations when only the marginal distributions for each bank are known (i.e., under the comonotonic copula). In every setting, we notice that the total minimal requirements are lower than the aggregate Nash allocations, which in turn are lower than the comonotonic copula case. In fact, the Nash allocations for each bank are lower than their counterpart under the comonotonic copula; this, however, does not hold between the minimal capital requirements and the Nash allocations. These systemic risk measures and capital allocations are plotted in Figure 1.
(a) Expectation risk measure
(b) Average Value-at-Risk
Figure 1: Section 4.3.1: Visualizations of the sensitive systemic risk measures, scenario-dependent systemic risk measures, and different allocation rules under different risk measures.
4.3.2 European Banking Data
Herein, we consider a larger system with $N=87$ banks. We calibrate this system to the 2011 EBA stress testing dataset as also undertaken in, e.g., Gandy and Veraart (2017); Chen et al. (2016); Feinstein et al. (2018). [^5] The exact calibration methodology is provided in Appendix B. In comparison to Section 4.3.1 above, in this example, we will consider a deterministic shock in which the initial external assets $X_{i}$ of bank $i\in[N]$ are as reported in the EBA dataset less 10% of its total exposures. [^6] Note that, as there is no randomness, the comonotonic copula of $X$ is itself. Due to the deterministic shock, every coherent risk measure provides the same results; therefore, we can completely characterize the set of acceptable allocations by the threshold $\gamma=97.5\%$ for the total repayment to society. For the purposes of this example, we will solely consider the sensitive systemic risk measure setting. (As this is a deterministic setting, the scenario-dependent systemic risk measure is the truncated halfspace defined by the total minimal capital requirements.) Recall from Corollary 4.5 that this system admits a unique Nash allocation.
Figure 2 displays the capital requirements assigned to each bank in the EBA dataset under the sensitive systemic risk measure setting. First, it is notable how the minimal capital requirements and the Nash allocation closely coincide for many banks in the system (Figure 2(b)). However, as observed in Figure 2(a), under the minimal capital requirements, there are a few institutions that “freeload” by having large negative capital requirements which need to be offset by other banks. As observed in Figure 2(a), some of the minimal capital allocations are of a larger order of magnitude than the Nash allocations. In comparison, the Nash allocations appear to be fairer in the sense that the requirements for all banks are closely aligned (but do differ due to bank-specific attributes). When we look at the total capital requirements, we find that the Nash allocations are less than €4.7 Million larger than the minimal acceptable capital allocation; the minimal total capital required is €728,316.4122 Million whereas the total Nash allocation is €728,321.1119 Million. Notably, this €4.7 Million is distributed throughout the $87$ institutions as only 4 banks have a lower capital requirement under the minimal capital allocation than under Nash.
(a) All capital allocations.
(b) Only banks with smaller capital requirements under the Nash allocation.
Figure 2: Section 4.3.2: Visualization of the Nash (x-axis) and minimal capital allocations (y-axis). The solid black line indicates the level where the Nash and minimal capital allocations coincide. Banks with larger capital requirements under the minimal capital allocations are blue circles; banks with larger capital requirements under the Nash allocations are red stars.
5 Conclusion
Within this work, we propose a novel and financially meaningful allocation mechanism for capital requirements. This approach, based on game theory in which institutions compete for lower capital requirements, allows regulators to assign capital requirements to banks while enforcing macro-prudential requirements. We find that the insensitive risk measures of Chen et al. (2013); Kromer et al. (2016) generally permit a unique Nash allocation rule, whereas the sensitive risk measures of Feinstein et al. (2017) may require additional regularity conditions for uniqueness. We investigate the Eisenberg and Noe (2001) clearing system for systemic risk measures and numerically illustrate the associated Nash allocations for the Eisenberg-Noe aggregation function.
While we focus on the existence and uniqueness of Nash allocation rules within this work, other mathematical and financial properties of these rules need to be investigated. For instance, we wish to consider if any of (lower semi-)continuity, monotonicity, translativity, or convexity can be recovered for a Nash allocation rule. We highlight that, as per Ararat and Feinstein (2024), at least one of these desirable properties cannot hold generally as interactions between firms is guaranteed by the construction.
Acknowledgments
The first author is partially supported by TÜBİTAK (Scientific & Technological Research Council of Turkey), Project No. 123F357.
Appendix A Nash Allocations with Rescue Funds
Within the main body of this text, we define Nash allocations so that each firm holds onto its own capital. However, as shown in, e.g., Rogers and Veraart (2013), there are cases in which rescuing another bank may lead to better outcomes for the rescuer as well as the rescued firm. With that in mind, we propose the following variation on Definition 3.5.
Definition A.1.
Let $(\Lambda_{i})_{i\in[N]}$ be a decomposition of $\Lambda$. A matrix-valued functional $r\colon L^{\infty}_{\Lambda}(\mathbb{R}^{N})\to\mathbb{R}^{N\times N}$ is called a $(\mathcal{A},(\Lambda_{i})_{i\in[N]})$-Nash allocation rule with rescue funds if, for every $X\in L^{\infty}_{\Lambda}(\mathbb{R}^{N})$ and $i\in[N]$, it holds
$r_{i\cdot}(X)\in\operatorname*{arg\,min}_{m_{i}\in\mathbb{R}^{N}}\left\{\sum_{j\in[N]}m_{ij}\;\left|\;\begin{array}[]{l}\Lambda_{i}(X,m_{i}+\sum_{j\neq i}r_{j\cdot}(X))\in\mathcal{A},\\ m_{ii}e_{i}+\sum_{j\neq i}r_{jj}(X)e_{j}\in D(X),\;m_{ij}\geq 0\,\forall j\neq i\end{array}\right.\right\}.$
In this case, $r$ is called acceptable if $\sum_{i\in[N]}r_{i\cdot}(X)\in R(X)$ for every $X\in L^{\infty}_{\Lambda}(\mathbb{R}^{N})$.
A Nash allocation rule with rescue fund generalizes the Nash allocation rules introduced within Definition 3.5. Specifically, this variation allows for firm $i$ to provide a rescue $m_{ij}\geq 0$ to firm $j$. Notably, these rescues must be nonnegative to restrict banks from stealing funds from other firms in the system. [^7] An additional feasibility condition is imposed which guarantees the finiteness of the aggregation function for each bank when neglecting the cash transfers. This condition is imposed rather than, for bank $i$, $m_{i}+\sum_{j\neq i}r_{j\cdot}(X)\in D(X)$ to preclude bank $i$ from using the provided transfers $r_{ji}(X)$ to guarantee that the system is feasible, i.e., each bank guarantees that the aggregation is well-defined even if there is a delay in the transfer payments. Overall, each institution is still attempting to minimize the total capital it is injecting into the system.
We conclude this discussion of the Nash allocation rules with rescue funds by considering when this generalization reduces to Definition 3.5. Specifically, we find that the self-preferential property guarantees that any Nash allocation rule with rescue funds will not include any transfer payments.
Proposition A.2.
Let $(\Lambda_{i})_{i\in[N]}$ be a self-preferential decomposition of $\Lambda$. Let $r\colon L^{\infty}_{\Lambda}(\mathbb{R}^{N})\to\mathbb{R}^{N\times N}$ define a $(\mathcal{A},(\Lambda_{i})_{i\in[N]})$-Nash allocation rule with rescue funds. Then, for every $X\in L^{\infty}_{\Lambda}(\mathbb{R}^{N})$ and $i,j\in[N]$ with $i\neq j$, it holds $r_{ij}(X)=0$. In particular, $(r_{11},\ldots,r_{NN})^{\mathsf{T}}$ is a $(\mathcal{A},(\Lambda_{i})_{i\in[N]})$-Nash allocation rule as defined in Definition 3.5.
Proof.
Let $L<1$ be the self-preferential constant associated with $(\Lambda_{i})_{i\in[N]}$. Let us fix a stress scenario $X\in L^{\infty}_{\Lambda}(\mathbb{R}^{N})$. Consider bank $i\in[N]$ and assume that $\sum_{j\neq i}r_{ij}(X)>0$. By the monotonicity and self-preferential property of the decomposition, we have
$\Lambda_{i}\left(X,r_{i\cdot}(X)+\sum_{j\neq i}r_{j\cdot}(X)\right)\leq\Lambda_{i}\left(X,\left(r_{ii}(X)+L\sum_{j\neq i}r_{ij}(X)\right)e_{i}+\sum_{j\neq i}r_{j\cdot}(X)\right).$
Notably, $r_{ii}(X)+L\sum_{j\neq i}r_{ij}(X)<\sum_{k\in[N]}r_{ik}(X)$, which contradicts the optimality of $r_{i\cdot}(X)$. ∎
Appendix B Calibration to European Banking Dataset
Within Section 4.3.2, we reconstruct the interbank liability matrix following the approach outlined in Gandy and Veraart (2017). In doing so, we exactly copy the hyperparameters utilized in Feinstein (2019).
First, we note that the 2011 EBA stress testing dataset provides the total assets $T_{i}$, capital $c_{i}$, and interbank liabilities $\sum_{j\in[N]}\bar{p}_{ij}$ for all banks $i\in[N]$ in the dataset. [^8] Due to technical issues with the calibration, we only consider $N=87$ of the $90$ institutions in the dataset. DE029, LU45, and SI058 are not included within this calibration. For simplicity, we set the total interbank assets to be equal to the total interbank liabilities, i.e., $\sum_{j\in[N]}\bar{p}_{ij}=\sum_{j\in[N]}\bar{p}_{ji}$ for every bank $i$ with slight perturbations as discussed within Gandy and Veraart (2017). For simplicity, we assume that the difference between the total assets ($T_{i}$) and the interbank assets ($\sum_{j\in[N]}\bar{p}_{ji}$) are the (initial) external assets ($x_{i}:=T_{i}-\sum_{j\in[N]}\bar{p}_{ji}$). Similarly, all liabilities that are not owed to other banks are external ($\bar{p}_{i0}:=T_{i}-\sum_{j\in[N]}\bar{p}_{ij}-c_{i}$).
With these parameters, we reconstruct the full nominal liabilities matrix, i.e., the exact values of $\bar{p}_{ij}$ for $i\neq j$, using the methodology of Gandy and Veraart (2017). Though a stochastic framework, within Section 4.3.2, we only consider a single realization of this liabilities matrix. For this reconstruction we consider hyperparameters $p=0.5$, $\text{thinning}=10^{4}$, $n_{\text{burn-in}}=10^{9}$, and $\lambda=\frac{pn(n-1)}{\sum_{i\in[N]}\sum_{j\in[N]}\bar{p}_{ij}}\approx 0.00122$.